### v1

# SWI: Speaking with Intent in Large Language Models

Abstract
Intent, typically clearly formulated and planned, functions as a cognitive framework for reasoning and problem-solving. This paper introduces the concept of Speaking with Intent (SWI) in large language models (LLMs), where the explicitly generated intent encapsulates the model’s underlying intention and provides high-level planning to guide subsequent analysis and communication. By emulating deliberate and purposeful thoughts in the human mind, SWI is hypothesized to enhance the reasoning capabilities and generation quality of LLMs. Extensive experiments on mathematical reasoning benchmarks consistently demonstrate the superiority of Speaking with Intent over Baseline (i.e., generation without explicit intent). Moreover, SWI outperforms answer-trigger prompting methods Chain-of-Thought and Plan-and-Solve and maintains competitive performance with the strong method ARR (Analyzing, Retrieving, and Reasoning). Additionally, the effectiveness and generalizability of SWI are solidified on reasoning-intensive question answering (QA) and text summarization benchmarks, where SWI brings consistent improvement to the Baseline generation. In text summarization, SWI-generated summaries exhibit greater accuracy, conciseness, and factual correctness, with fewer hallucinations. Furthermore, human evaluations verify the coherence, effectiveness, and interpretability of the intent produced by SWI. This proof-of-concept study creates a novel avenue for enhancing LLMs’ reasoning abilities with cognitive notions. [^1] Appendix A Reproducibility Statement. Source code: https://github.com/YuweiYin/SWI
1 Introduction
Large language models (LLMs) (Zhao et al., 2023; Min et al., 2023; Minaee et al., 2024) have revolutionized Natural Language Processing (NLP) with their excellent generative capabilities (Hurst et al., 2024; Anthropic, 2024; Team et al., 2024). Many recent benchmarks focus on evaluating complex reasoning in mathematics problems (Cobbe et al., 2021; Hendrycks et al., 2021b; Lightman et al., 2024; Vendrow et al., 2025) and multidisciplinary questions (Hendrycks et al., 2021a; Srivastava et al., 2023; Suzgun et al., 2023; Wang et al., 2024e), presenting significant challenges for LLMs. Enhancing LLM reasoning is vital for their ongoing development (Huang & Chang, 2022; Qiao et al., 2023; Patil, 2025).
Recent advancements have introduced various approaches to improve LLM reasoning (Yu et al., 2024; Cheng et al., 2025; Liu et al., 2025a; Xu et al., 2025). Chain-of-Thought (CoT) prompting enhances reasoning (Li et al., 2024b) by providing step-by-step solutions in few-shot exemplars (Wei et al., 2022) or explicitly instructing the model to “think step by step” (Kojima et al., 2022). Beyond CoT variants (Yao et al., 2023a; Wang et al., 2023; Yasunaga et al., 2024), Yin & Carenini (2025) propose Analyzing, Retrieving, and Reasoning (ARR), an answer-trigger prompting method that systematically guides LLMs in question answering and outperforms CoT prompting.
Inspired by ARR (Yin & Carenini, 2025), where intent analysis is identified as its most effective component, we further investigate the powerful potential that intent can unleash in enhancing LLM reasoning. Intent, the goal-oriented intention in our mind (Adams, 1986; Mele, 1989; Mele & Moser, 1994), serves as a guiding framework for problem-solving. As illustrated in Figure 1(a), human reasoning (System 2 (Kahneman, 2011)) typically follows a structured thinking loop where intent directs problem analysis and logical reasoning. Hence, we hypothesize that enabling LLMs to explicitly speak with their own intent—rather than merely analyzing the intent of questions—can replicate this meta-cognitive planning process, thereby improving the reasoning ability and generation quality of LLM.
Figure 1: SWI Overview. (a) The intent, functioning as meta-thought and planning, guides the analysis with reasoning to answer the question. (b) A concrete example generated by our method: Speaking with Intent (SWI) in large language models (LLMs). (c) The performance improvement on mathematical reasoning and multiple-choice QA tasks brought by SWI.
This work introduces Speaking with Intent (SWI), requiring LLMs to articulate their own intent as a planning mechanism during generation. Due to the autoregressive nature of LLMs (Radford et al., 2019) and the attention mechanism (Vaswani et al., 2017), explicitly stated intent provides high-level guidance for subsequent analysis and reasoning. Specifically, we implement SWI using instruction-following LLMs (Ouyang et al., 2022; Rafailov et al., 2023), incorporating system and user prompts for the model (Dubey et al., 2024) to generate intent. As shown in Figure 1(b), each analytical step (in blue) in solving a math problem is guided by a preceding intent statement (in orange), which is a freely generated text instead of a predefined class as in traditional intent modeling (Weld et al., 2022).
To comprehensively evaluate the effectiveness and generalizability of the proposed SWI method, we conduct experiments across three diverse task categories: mathematical reasoning, multiple-choice QA, and text summarization. As shown in Figure 1(c), LLM speaking with intent (SWI) consistently outperforms Baseline, which generates responses without explicit intent. Furthermore, SWI surpasses answer-trigger prompting methods such as Chain-of-Thought (CoT) (Kojima et al., 2022) and Plan-and-Solve (PS) (Wang et al., 2023). Compared with the strong method ARR (Analyzing, Retrieving, and Reasoning) (Yin & Carenini, 2025), SWI maintains competitive performance on mathematical reasoning tasks. For text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations (Ji et al., 2023; Li et al., 2024a) in the output.
Additionally, we perform human evaluations to assess the coherence, effectiveness, and interpretability of the intent generated by our SWI method. Evaluators largely agree on the quality of the generated intent, particularly for mathematical reasoning tasks. The evaluation results confirm that Speaking with Intent in LLMs enhances task performance and output explainability. The key contributions of this work are as follows:
  • 1.
    This paper introduces Speaking with Intent (SWI) in LLMs, where the generated intent effectively guides problem analysis and logical reasoning, improving performance across various benchmarks. The source code will be publicly available.
  • 2.
    Extensive experiments across diverse reasoning and generation tasks, including mathematical reasoning, multiple-choice QA, and text summarization, demonstrate that SWI consistently outperforms the Baseline and surpasses established answer-trigger prompting methods such as Chain-of-Thought.
  • 3.
    Human evaluations validate the coherence, effectiveness, and interpretability of the intent generated by SWI. Rather than intent classification, our evaluation practice provides standards for assessing freely generated intent statements.
2 Related Work
LLM Reasoning.
Recent LLM research increasingly focuses on enhancing reasoning abilities (Sun et al., 2023; Patil, 2025; Cheng et al., 2025; Liu et al., 2025a). Chain-of-Thought (CoT) prompting facilitates problem-solving by prompting LLMs to generate intermediate reasoning steps (Kojima et al., 2022; Wei et al., 2022; Li et al., 2024b; Yeo et al., 2025). Building on CoT, various reasoning techniques have emerged (Xu et al., 2025). Among them, a recent research ARR (Yin & Carenini, 2025) consistently outperforms CoT on multiple QA tasks, where intent analysis of questions is its most effective component. Inspired by ARR, we enable LLMs to articulate their intent, using it to guide subsequent analysis and reasoning for improved task performance. Results show that Speaking with Intent (SWI) surpasses CoT and ARR on multiple benchmarks. As an intent-driven approach, SWI introduces a novel framework for advancing LLM reasoning and cognitive capabilities.
LLM Planning.
Planning plays a critical role in advancing artificial intelligence (Wang et al., 2024c), particularly in developing LLM agents in real and complex environments (Qiao et al., 2024; Mialon et al., 2024; Xu et al., 2024; Drouin et al., 2024; Boisvert et al., 2024; Xie et al., 2024a; b). Aside from improving the tool-use capabilities (Wang et al., 2024f; Wu et al., 2024; Huang et al., 2024; Wang et al., 2024a; Liu et al., 2025b), recent LLM planning methods enhance reasoning abilities (Huang et al., 2022; Yao et al., 2023b; Hao et al., 2023; Valmeekam et al., 2023; Jiao et al., 2024) by generating problem-solving plans explicitly (Wang et al., 2023; Li & Zhang, 2024; Lyu et al., 2024) or implicitly (Goyal et al., 2024; Wang et al., 2024d; Pfau et al., 2024; Chen et al., 2024). In this work, we require LLMs to explicitly speak with their intent in an iterative and dynamic manner, as shown in Figure 1(a), instead of generating all plans first and then solving the task (Wang et al., 2023; Li & Zhang, 2024). The existing iterative planning framework (Lyu et al., 2024) focuses on addressing information retrieval issues in knowledge-intensive long-form generation tasks, while our SWI method generates intents that function as strategic planning to guide problem analysis and logical reasoning.
Intent-related Research.
Intent Detection (ID) and New Intent Discovery (NID) (Kumar et al., 2022; Liang et al., 2024; Zhang et al., 2024a; Tang et al., 2024; Zhang et al., 2024c; Qian et al., 2024; Yin et al., 2025), which classify utterances into known or novel intent categories, are longstanding challenges in natural language understanding (Larson et al., 2019; Casanueva et al., 2020; Zhang et al., 2021; Weld et al., 2022). Typically, these tasks are approached as classification problems (Wang et al., 2024b; Yoon et al., 2024; Zhang et al., 2024b; Sakurai & Miyao, 2024), where models assign sentences to predefined intent classes. In contrast, Speaking with Intent (SWI) generates intent as free-form text rather than fixed categories, enhancing flexibility and fluency. SWI naturally integrates intent statements as planning into the reasoning process, providing contextual guidance for subsequent analysis.
3 Speaking with Intent
This section presents the problem-solving workflow of LLMs and introduces Speaking with Intent (SWI), enabling LLMs to explicitly articulate their intent for planning and reasoning.
3.1 Problem-solving Workflow using LLMs
Let $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ be a dataset, where $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n}\}$ is the input information (questions), $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$ is the corresponding references (correct answers), and $n$ is the number of instances in $\mathcal{D}$. For mathematical reasoning datasets, $X_{i}$ is the math problem, and $y_{i}$ is the correct answer (usually an integer number). For multiple-choice QA datasets, $X_{i}$ contains the question and options, and $y_{i}$ is the answer label such as (A)/(B)/(C) or Yes/No. For text summarization datasets, $X_{i}$ is the source article, and $y_{i}$ is the reference summary.
In this work, we employ instruction-following LLMs (aka Chat LLMs) $\mathcal{M}$ for experiments and apply the chat template with the system prompt $P_{s}$ and user prompt $P_{u}$. The system prompt specifies the general behavior of the model (assistant), and the user prompt poses questions to the model. Therefore, the generated output $\hat{y}_{i}$ is obtained by
$\hat{y}_{i}=\mathcal{M}(P_{s},P_{u},X_{i};\Theta,\zeta),$ (1)
[…]
3.2 Baseline vs. Speaking with Intent
In this work, we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts $P_{s}$ and restating the SWI requirement in the user prompt $P_{u}$. Figure 2 presents the system prompt and user prompt of Baseline generation ($P_{s}^{\text{base}}$ and $P_{u}^{\text{base}}$) and our SWI method ($P_{s}^{\text{swi}}$ and $P_{u}^{\text{swi}}$).
Figure 2: LLM Prompts. The system and user prompts of Baseline and our SWI method.
3.3 Result Evaluation
Lastly, we extract the final answer (after “Final Answer:” in $\hat{y}_{i}$), denoted as $\tilde{y}_{i}$, and compute the overall performance of $\mathcal{M}$ on the dataset $\mathcal{D}$ by
$s=\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(y_{i},\tilde{y}_{i}),$ (2)
where the score function $\mathcal{S}(\cdot,\cdot)$ returns a value in the range of $[0,1]$. Different tasks adopt different score functions to evaluate the model performance.
For mathematical reasoning tasks, we first extract numbers in $\tilde{y}_{i}$ and apply text normalization to both $\tilde{y}_{i}$ and the reference $y_{i}$, and then conduct exact match to check if the generated answer $\tilde{y}_{i}$ is correct.
For multiple-choice QA tasks, we adopt the Option Selection metric introduced by Yin & Carenini (2025), which evaluates the LLM perplexity of different option concatenations and selects the one with lowest perplexity as the model’s choice.
To evaluate the quality of summaries, we apply the standard ROUGE (Lin, 2004) as the automatic evaluation metric $\mathcal{S}$ and complement it with a more sophisticated fact-checking analysis as described in Section 5.2.
4 Main Experiments on Mathematical Reasoning
This section presents the main experiments to verify the effectiveness of SWI on mathematical reasoning. To ensure reproducibility, we fixed all random seeds, set the generation temperature to zero, and conducted all experiments twice, yielding reproducible results. [^2] Please refer to Appendix A for the reproducibility statement and Appendix B for dataset details.
4.1 Models
In all experiments, we employ LLaMA3-8B-Instruct (Dubey et al., 2024) as the model $\mathcal{M}$ for generation and evaluation. It is an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters. We use the model checkpoint and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020).
4.2 Datasets
The proposed SWI method aims to enhance LLM reasoning, so we conduct extensive experiments on various mathematical reasoning benchmarks, ranging from grade-school to competition-level problems. In the math reasoning task, the model is asked to solve the given math problem and present the final answer. We consider grade-school math problems including GSM8K (Cobbe et al., 2021), GSM8K-Platinum (Vendrow et al., 2025), and MATH500 (Lightman et al., 2024), as well as three competition-level math benchmarks: AMC23, AIME24, and AIME25, which have 40, 30, and 30 math problems, respectively. Despite the smaller scale, these math problems are much more challenging than those in GSM8K and MATH500. The dataset statistics are presented in Table 1.
[…]
Mathematical Reasoning GSM8K (Cobbe et al., 2021) Test 1,319 124
GSM8K-P (Vendrow et al., 2025) Test 1,209 124
MATH500 (Lightman et al., 2024) Test 500 134
[…]
4.3 Experimental Comparison
The main comparison is LLM generation with intent (SWI) or without intent (Baseline), and the effectiveness of SWI is verified if the former outperforms the latter. For mathematical reasoning and multiple-choice QA tasks, we also compare SWI with answer-trigger prompting methods CoT (Kojima et al., 2022) and ARR (Yin & Carenini, 2025), as well as previous LLM planning method Plan-and-Solve (PS) prompting (Wang et al., 2023).
CoT aims to elicit LLM reasoning using the answer-trigger prompt $\Phi_{i}^{\text{CoT}}$ as “Let’s think step by step”, while ARR, with a more systematic design, utilizes an enhanced $\Phi_{i}^{\text{ARR}}$ as “Let’s analyze the intent of the question, find relevant information, and answer the question with step-by-step reasoning.” Finally, PS applies the following prompt $\Phi_{i}^{\text{PS}}$ to construct plans before problem-solving: “Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan and solve the problem step by step.” With answer-trigger prompts $\Phi_{i}$, the generation process (Eq. 1) is given by
$\hat{y}_{i}=\mathcal{M}(P_{s},P_{u},X_{i},\Phi_{i};\Theta,\zeta),$ (3)
where $P_{s}$ is the system prompt, $P_{u}$ is the user prompt, and $X_{i}$ is the task input. $\Theta$ and $\zeta$ are the parameters and hyper-parameters of the LLM $\mathcal{M}$, respectively.
4.4 Experimental Results
SWI vs. Baseline.
As shown in Table 2, SWI improves Baseline (generation without intent) performance by an average of +3.34 points, showing its effectiveness in enhancing LLM’s problem analysis and mathematical reasoning abilities.
Method GSM8K GSM8K-P MATH500 Competition Avg.
Baseline 78.47 81.14 36.80 10.56 51.74
SWI (ours) 81.20 84.20 41.60 13.33 55.08
Table 2: Mathematical Reasoning Results. The exact matching scores (%) of LLaMA-3-8B-Instruct on multiple mathematical reasoning datasets with or without Speaking with Intent.
SWI vs. Answer-Trigger Prompting Methods.
Table 3 presents the results of our SWI method and three answer-trigger prompting methods: CoT, PS, and ARR. We observe that SWI significantly outperforms PS, suggesting that the dynamic generation of intents in SWI provides more effective analytical guidance than the static construction of multiple plans at the outset, as in PS. Furthermore, compared to the reasoning-eliciting methods CoT and ARR, SWI demonstrates superior performance over CoT and achieves competitive results with ARR. Notably, SWI surpasses ARR in solving competition-level mathematics problems, which are more challenging and thus indicate the potential of SWI.
Method GSM8K GSM8K-P MATH500 Competition Avg.
CoT (Kojima et al., 2022) 79.30 82.22 42.60 7.78 52.98
PS (Wang et al., 2023) 73.54 76.43 39.20 8.61 49.45
ARR (Yin & Carenini, 2025) 81.96 85.36 44.40 10.28 55.50
SWI (ours) 81.20 84.20 41.60 13.33 55.08
Table 3: Mathematical Reasoning Results. The exact matching scores (%) of LLaMA-3-8B-Instruct on multiple mathematical reasoning datasets using our SWI method or answer-trigger prompting methods. The underlined values indicate the second-best scores.
5 Generalizability of Speaking with Intent
The main experiments have verified the effectiveness of Speaking with Intent, this section further studies the generalizability of the proposed SWI method on different categories of tasks: multiple-choice QA and text summarization.
5.1 Multiple-choice Question Answering
Many challenging benchmarks are designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question.
5.1.1 QA Datasets
Our SWI method is evaluated on various QA datasets involving reading comprehension (Liu et al., 2020), commonsense reasoning (Talmor et al., 2019; Sap et al., 2019), world knowledge (Mihaylov et al., 2018), and reasoning-intensive multitask understanding tasks with multiple subtasks covering diverse topics (Clark et al., 2018; Suzgun et al., 2023; Hendrycks et al., 2021a; Wang et al., 2024e). The dataset statistics are presented in Table 4.
[…]
Reading LogiQA (Liu et al., 2020) Test 651 185
Commonsense CSQA (Talmor et al., 2019) Valid 1,221 127
SIQA (Sap et al., 2019) Valid 1,954 117
World Knowledge OBQA (Mihaylov et al., 2018) Test 500 127
ARC (Clark et al., 2018) Test 3,548 144
Multitask Understanding BBH (Suzgun et al., 2023) Test 5,511 203
MMLU (Hendrycks et al., 2021a) Test 13,842 193
MMLU-Pro (Wang et al., 2024e) Test 12,032 270
[…]
5.1.2 QA Results
Table 5 shows that our SWI method consistently improves the Baseline (generation without intent) by a large margin, i.e., +10.08 points on average. The results demonstrate the efficacy of Speaking with Intent in reasoning-intensive question answering.
Method Reading Commonsense World Knowledge Multitask Understanding Avg.
LogiQA CSQA SIQA OBQA ARC BBH MMLU MMLU-Pro
Baseline 20.74 66.91 29.58 70.60 76.44 56.65 52.40 39.27 51.57
SWI (ours) 33.03 72.73 46.37 81.60 87.82 64.29 62.19 45.16 61.65
Table 5: Multiple-choice QA Results. The accuracy scores (%) of LLaMA3-8B-Instruct on various multiple-choice QA datasets with or without Speaking with Intent.
5.2 Text Summarization
The results on math and QA benchmarks have solidified the efficacy of SWI on reasoning-intense tasks. Beyond reasoning tasks, we hypothesize that Speaking with Intent benefits natural language generation tasks like summarization by more explicitly analyzing the source document point by point and better planning the generation of the final summary.
5.2.1 Summarization Datasets
To verify the hypothesis, we test the effect of SWI on various text summarization datasets. CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017), Extreme summarization (XSum) (Narayan et al., 2018), and XL-Sum (Hasan et al., 2021) contain diverse article-summary pairs from news articles. SAMSum (Gliwa et al., 2019) and DialogSum (Chen et al., 2021) are dialogue summarization datasets. WikiLingua (Ladhak et al., 2020) extracts article and summary pairs from wikiHow. The dataset statistics are presented in Table 6.
[…]
News Article CDM (Hermann et al., 2015) Test 11,490 920
XSum (Narayan et al., 2018) Test 11,334 542
XL-Sum (Hasan et al., 2021) Test 11,535 613
Dialogue SAMSum (Gliwa et al., 2019) Test 819 209
DialogSum (Chen et al., 2021) Test 1,500 263
Wiki Article WikiLingua (Ladhak et al., 2020) Test 3,000 525
[…]
5.2.2 Summarization Results
Automatic Evaluation of Summaries.
We evaluate the quality of summaries using the ROUGE score (Lin, 2004), which counts the overlaps of the generated summaries and reference summaries. Specifically, we average the ROUGE-1 (unigrams), ROUGE-2 (bigrams), ROUGE-L (longest common subsequences), and ROUGE-LSum (sentence-level ROUGE-L) scores as the final ROUGE score. As shown in Table 7, our SWI method consistently surpasses the Baseline on ROUGE scores, confirming its effectiveness in enhancing the quality of text summaries.
Method News Article Dialogue Wiki Article Avg.
CDM XSum XL-Sum SAMSum DialogSum WikiLingua
Baseline 23.38 11.90 11.29 24.14 16.92 15.01 17.11
SWI (ours) 24.00 13.82 13.50 24.37 20.64 17.86 19.03
Table 7: Text Summarization Results. The ROUGE scores (%) of the LLaMA3-8B-Instruct model on various text summarization datasets with or without Speaking with Intent (SWI).
5.2.3 Fact Checking of Summaries
LLMs frequently generate hallucinated content (Ji et al., 2023; Li et al., 2024a), which can not be detected by lexical metrics like ROUGE. To assess this issue, we adopt a more semantically sophisticated fact-checking metric proposed by Hwang et al. (2025), which quantifies factual consistency by calibrating the extent of fabricated statements (low precision) and omitted factual information (low recall). Specifically, we use GPT-4o-mini (Hurst et al., 2024) to decompose both generated and reference summaries into atomic fact sets and measure their coverage to quantify factual consistency.
We evaluate $100$ samples from each summarization dataset using this fact-checking metric, with results presented in Table 8. Speaking with Intent (SWI) generates more concise and accurate summaries, improving precision, whereas baseline summaries tend to be more lengthy and verbose, resulting in higher recall scores. Overall, SWI consistently outperforms the Baseline generation in terms of F1 score.
Dataset Method Precision Recall F1 Dataset Method Precision Recall F1
CDM Baseline 26.06 76.28 36.37 SAMSum Baseline 39.95 75.92 46.90
SWI 34.22 55.89 37.79 SWI 47.82 69.88 50.88
XSum Baseline 11.06 48.38 15.15 DialogSum Baseline 23.99 57.08 29.55
SWI 14.77 37.30 16.29 SWI 34.92 45.19 31.20
XL-Sum Baseline 8.96 61.88 13.79 WikiLingua Baseline 23.33 65.55 30.63
SWI 12.96 46.72 16.51 SWI 32.40 54.98 35.78
Table 8: Summarization evaluation results. We compare the atomic facts drawn from the generated and reference summaries, and compute recall, precision, and F1 scores (%).
5.2.4 Case Study
To provide more insights into the benefits of Speaking with Intent, we conduct a case study on the CDM summarization dataset (Hermann et al., 2015; See et al., 2017). Figure 3 showcases the SWI outputs, where the generated intent is well formulated and articulated, guiding a point-by-point summarization process that leads to a final summary that is accurate, concise, and abstractive, while effectively capturing the key information of the source article. This probably leads to the high Precision and F1 scores of SWI in Table 8, suggesting the validity of Speaking with Intent in text generation tasks.
Figure 3: Case Study. Task input and SWI output on text summarization task.
6 Intent Quality Evaluation
Although we have shown that Speaking with Intent boosts performance across a broad range of tasks, verifying the quality of generated intents is also significant. Thus, we hire human evaluators to assess the quality of generated intent across three criteria: coherence, effectiveness, and interpretability. Coherence measures how well the intent guides analysis and reasoning, effectiveness evaluates its contribution to problem-solving, and interpretability assesses its role in enhancing user understanding of the generated content. [^3] Please refer to Appendix C for more details on human evaluation. For each instance, human evaluators are provided with evaluation instructions, task input (e.g., the math problem, question with options, or source article), SWI-generated output, and assessment check boxes. They are then asked to evaluate the following aspects:
  • Coherence: In general, does the analysis align coherently with the intent statements?
  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for solving the problem?
  • Interpretability: Do you think providing the intent can help you better understand the reasoning process than not providing it?
Evaluation scores range from 1 (Bad), 2 (Fair), to 3 (Good). Agreement ratios are calculated as follows: 1 if all three evaluators agree, 0.5 if two agree, and 0 if all scores differ. As shown in Table 9, human evaluation scores for all aspects across datasets exceed 2, indicating that the generated intent is generally well-regarded. Notably, math reasoning tasks achieve scores near 3 with agreement ratios approaching 100%, demonstrating that SWI-generated intent is particularly coherent, effective, and interpretable in mathematical problem-solving.
The relatively low scores observed in QA tasks may be attributed to the lack of multi-step guidance: the number of generated intents in QA tasks is often 1 or 2, while that in math and summarization is usually at least three. This finding indicates the advantages of multi-round iterative intents, although SWI can boost task performance even with few intents generated.
Task Dataset Coherence Effectiveness Interpretability
Score Agree Score Agree Score Agree
Math Reasoning GSM8K 2.90 85% 2.97 95% 2.97 95%
MATH500 2.87 80% 2.87 80% 2.83 80%
Multiple-choice QA BBH 2.37 55% 2.37 50% 2.33 45%
MMLU 2.67 75% 2.53 55% 2.37 45%
Summarization CDM 2.83 80% 2.77 70% 2.83 75%
XSum 2.70 70% 2.60 65% 2.57 65%
Table 9: Intent Quality Evaluation by Humans. The score ranges from 1 (Bad) to 3 (Good).
7 Conclusion
In this work, we introduce Speaking with Intent (SWI) in LLMs, where the generated intent (as high-level planning) guides subsequent analysis, improving the reasoning and generation abilities. Extensive experiments across mathematical reasoning, multiple-choice QA, and text summarization benchmarks consistently show the benefits of SWI over the Baseline method that generates without explicit intent. In addition, SWI outperforms answer-trigger prompting methods like Chain-of-Thought on mathematical reasoning benchmarks. In text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations. Furthermore, human evaluations solidify the coherence, effectiveness, and interpretability of LLM-generated intent. This study opens a new avenue for enhancing LLM reasoning abilities and beyond.
Impact Statement.
As intent is a fundamental aspect of natural language processing, empowering, eliciting, and enhancing the intent understanding and generation abilities can potentially drive AI systems (including multimodal models) to the next level. Moreover, Speaking with Intent can also be applied to various domains beyond NLP, such as healthcare, law, and finance. These applications are cost-sensitive, so explicitly showing the intent of AI models will help with the transparency and interpretability of critical decision-making.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), Vector Institute for AI, and Accelerate Foundation Models Research Program Award from Microsoft. This research was supported in part by the computational resources and services provided by Advanced Research Computing at the University of British Columbia and the Digital Research Alliance of Canada (alliancecan.ca). We would also like to thank Vered Shwartz for constructive feedback on conducting evaluations.
Appendix A Reproducibility Statement
A.1 LLM Generation
For all experiments in § 4 and § 5, we employ LLaMA3-8B-Instruct (Dubey et al., 2024), an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters, and use the model checkpoint and tokenizer [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct provided by Hugging Face Transformers (Wolf et al., 2020). Only in the fact-checking evaluation for summaries (§ 5.2.3), we adopt the proprietary model GPT-4o-mini (Hurst et al., 2024) to decompose the generated summary and reference summary into two sets of atomic facts.
For each running session, the experiments are conducted on a single NVIDIA V100 GPU with 32GB memory. To avoid out-of-memory issues, all the models are loaded in a half-precision (float16) mode, and the generation batch size is $1$. The input sequence is not truncated since we do not want to lose the context information, but we set the maximum number of newly generated tokens as $4096$ during generation.
To guarantee reproducibility, we fixed the seeds as 42 for all random modules, set the LLM generation temperature as 0 for deterministic generation without sampling, and ran all experiments twice, obtaining reproducible generation results and evaluation scores. The source code is available on GitHub: https://github.com/YuweiYin/SWI
A.2 Experimental Cost
In the generation stage, the total computational cost is approximately 1,500 GPU hours on NVIDIA V100 clusters (about 58 days). In the evaluation stage, only the option selection process for multiple-choice QA tasks requires GPU usage (Yin & Carenini, 2025), and the overall running time is roughly 100 hours on V100 clusters. In the fact-checking experiments (§ 5.2.3), the expense for GPT-4o-mini API calls is below US\$10.
Appendix B Tasks and Datasets
To comprehensively study the effectiveness and generalizability of the proposed SWI method, we consider three different categories of tasks in our experiments: mathematical reasoning, multiple-choice QA, and text summarization. Table 10 presents the dataset sources. Please note that the URLs may be subject to change by the dataset providers.
Task Dataset URL
Mathematical Reasoning GSM8K (Cobbe et al., 2021) Link
GSM8K-P (Vendrow et al., 2025) Link
MATH500 (Lightman et al., 2024) Link
AMC23 (American Mathematics Competitions) Link
AIME24 (American Invitational Math Examination) Link
AIME25 (American Invitational Math Examination) Link
Multiple-choice QA LogiQA (Liu et al., 2020) Link
CSQA (Talmor et al., 2019) Link
SIQA (Sap et al., 2019) Link
OBQA (Mihaylov et al., 2018) Link
ARC (Clark et al., 2018) Link
BBH (Suzgun et al., 2023) Link
MMLU (Hendrycks et al., 2021a) Link
MMLU-Pro (Wang et al., 2024e) Link
Text Summarization CDM (Hermann et al., 2015) Link
XSum (Narayan et al., 2018) Link
XL-Sum (Hasan et al., 2021) Link
SAMSum (Gliwa et al., 2019) Link
DialogSum (Chen et al., 2021) Link
WikiLingua (Ladhak et al., 2020) Link
Table 10: Dataset sources.
B.1 Mathematical Reasoning
In the mathematical reasoning task, the model is asked to solve the given math problem and present the final answer. We consider the following math datasets.
GSM8K.
Grade School Math 8K (GSM8K) (Cobbe et al., 2021) is a dataset of 8.5K high-quality linguistically diverse grade school math word problems. The dataset was created to support the task of question answering on basic mathematical problems that require multi-step reasoning.
GSM8K-P.
GSM8K-Platinum (Vendrow et al., 2025) is a revised version of the full test set of GSM8K (Cobbe et al., 2021). It revises the labels of mislabeled examples and removes any question that is determined to be poorly written (mostly due to ambiguity in the problem statement), providing a more accurate assessment of mathematical reasoning capabilities.
MATH500.
MATH500 (Lightman et al., 2024) consists of 500 questions from the MATH test set (Hendrycks et al., 2021b). The distribution of difficulty levels and subjects in MATH500 is representative of the MATH test set as a whole.
Competition.
We consider three competition-level math benchmarks: AMC23, AIME24, and AIME25, drawn from American Mathematics Competitions and American Invitational Mathematics Examination. AMC23, AIME24, and AIME25 have 40, 30, and 30 math problems, respectively. Despite the smaller scale, these math problems are much more challenging than those in GSM8K and MATH500.
B.2 Multiple-choice QA
Since many challenging benchmarks are designed as multiple-choice question-answering tasks, our SWI method is also evaluated on various reasoning-intense QA datasets, where the model is asked to select the most appropriate one from the given options to answer the question.
LogiQA.
LogiQA Liu et al. (2020) is a reading comprehension dataset that requires the model to have logical reasoning for question-answering.
CSQA.
CommonsenseQA (CSQA) (Talmor et al., 2019) examines the model on commonsense question-answering problems constructed using information from ConceptNet Speer et al. (2017).
SIQA.
SocialIQA (SIQA) (Sap et al., 2019) is a large-scale QA benchmark for commonsense reasoning about social situations, which probes emotional and social intelligence in everyday situations.
OBQA.
OpenBookQA (OBQA) (Mihaylov et al., 2018) asks the model to answer the question based on the given elementary-level science facts and broad commonsense knowledge.
ARC.
AI2 Reasoning Challenge (ARC) (Clark et al., 2018) contains grade-school science questions. It is divided into a Challenge and an Easy set, where the former contains only questions answered incorrectly by both a retrieval-based algorithm and a word co-occurrence algorithm.
BBH.
BIG-Bench Hard (BBH) (Suzgun et al., 2023) is a suite of challenging tasks filtered from BIG-Bench (Srivastava et al., 2023). Solving these problems often requires multi-step reasoning. In this work, $4$ (out of $27$) subtasks in BBH (word_sorting, object_counting, dyck_languages, and multistep_arithmetic_two) are discarded as they are not multiple-choice QA tasks.
MMLU.
MMLU (Hendrycks et al., 2021a) comprehensively measures the multitask language understanding abilities on $57$ subtasks including elementary mathematics, history, computer science, and more.
MMLU-Pro.
MMLU-Pro (Wang et al., 2024e) extends the mostly knowledge-driven MMLU benchmark by integrating more challenging, reasoning-focused questions and expanding the choice set from four to ten options.
B.3 Text Summarization
Beyond math reasoning and multiple-choice QA, we hypothesize that Speaking with Intent benefits natural language generation tasks like summarization, where the generated intent can guide the model in summarizing the source article point by point. Hence, we test the effect of SWI on the following text summarization datasets (English only).
CNN/DailyMail.
CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017) contains over 300K unique news articles written by journalists at CNN and the Daily Mail.
XSum.
Extreme summarization (XSum) (Narayan et al., 2018) is a real-world, large-scale dataset consisting of online articles from the British Broadcasting Corporation (BBC).
XL-Sum.
XL-Sum (Hasan et al., 2021) is a comprehensive and diverse dataset comprising 1.35 million professionally annotated article-summary pairs from BBC News. The dataset covers 45 languages and we only use the English subset.
SAMSum.
SAMSum (Gliwa et al., 2019) contains about 16K messenger-like conversations with summaries. Linguists were asked to create conversations similar to those they write on a daily basis, reflecting the proportion of topics of their real-life messenger conversations.
DialogSum.
DialogSum (Chen et al., 2021) is a large-scale dialogue summarization dataset, consisting of 13,460 dialogues with corresponding manually labeled summaries and topics.
WikiLingua.
WikiLingua (Ladhak et al., 2020) is a large-scale, multilingual dataset for evaluating cross-lingual abstractive summarization systems. It extracts article and summary pairs in 18 languages from wikiHow, and we only use the English subset.
Appendix C Human Evaluation Details
Participant Requirements.
We hire human evaluators from the cloud-sourcing platform CloudResearch [^5] https://www.cloudresearch.com/ to conduct human evaluation on the quality of the generated intent: coherence, effectiveness, and interpretability. To ensure the annotation quality, we apply several requirements to select qualified human evaluators, as shown in Table 11.
Type Requirements
Native Language English
Country of Residence Australia, Canada, Ireland, New Zealand, UK, US
Education Undergraduate student, Graduate student
Reputation Approved Projects Count: $\geq$ 1,000
Approval Rating: $\geq$ 90%
Table 11: The requirements for human evaluators.
Evaluation Tasks.
For each task category, we select two datasets: GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2024) for mathematical reasoning, BBH (Suzgun et al., 2023) and MMLU (Hendrycks et al., 2021a) for multiple-choice QA, and CDM (Hermann et al., 2015; See et al., 2017) and XSum (Narayan et al., 2018) for text summarization. We randomly sample 12 instances per dataset and divide them into two batches of six. Each batch includes a dummy instance with deliberately reversed intents to ensure evaluators are actively engaged rather than randomly selecting responses. Evaluator submissions are accepted or rejected based on completion time and performance on the dummy instance.
For each instance, human evaluators are provided with evaluation instructions, task input (e.g., the math problem, question with options, or source article), SWI-generated output, and assessment check boxes. They are then asked to evaluate the following aspects:
  • Coherence: In general, does the analysis align coherently with the intent statements?
  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for solving the problem?
  • Interpretability: Do you think providing the intent can help you better understand the reasoning process than not providing it?
Each batch is assessed by three different human evaluators, with each person uniquely assigned to only one batch. Evaluation scores range from 1 (Bad), 2 (Fair), to 3 (Good). Agreement ratios are calculated as follows: 1 if all three evaluators agree, 0.5 if two agree, and 0 if all scores differ.
Human Evaluation Quality.
As mentioned above, we decided to accept or reject the evaluator’s submission based on the task completion time and the results on the dummy instance that is deliberately modified to have a lower coherence. As a result, about 60% of the evaluators still rated the dummy instance as good coherence, meaning they failed the dummy test and potentially did not fully focus on the evaluation process, which poses a general caveat to the quality of cloud-sourcing annotations. Overall, we rejected about 10% of submissions that both failed the dummy test and took an unreasonably short time to complete the annotation. After rejecting them, we hired other evaluators until the intent quality evaluation was finished.
Human Evaluation Cost.
The pay rate for each human evaluator is US\$10 per hour, completing a batch of 6 instances takes an evaluator 10-15 minutes on average, and the total cost of the intent quality evaluation is about US\$120.

### v3

@@
 Abstract
-Intent, typically clearly formulated and planned, functions as a cognitive framework for reasoning and problem-solving.
-This paper introduces the concept of Speaking with Intent (SWI) in large language models (LLMs), where the explicitly generated intent encapsulates the model’s underlying intention and provides high-level planning to guide subsequent analysis and communication.
+Intent, typically clearly formulated and planned, functions as a cognitive framework for communication and problem-solving.
+This paper introduces the concept of Speaking with Intent (SWI) in large language models (LLMs), where the explicitly generated intent encapsulates the model’s underlying intention and provides high-level planning to guide subsequent analysis and action.
 By emulating deliberate and purposeful thoughts in the human mind, SWI is hypothesized to enhance the reasoning capabilities and generation quality of LLMs.
-Extensive experiments on mathematical reasoning benchmarks consistently demonstrate the superiority of Speaking with Intent over Baseline (i.e., generation without explicit intent).
-Moreover, SWI outperforms answer-trigger prompting methods Chain-of-Thought and Plan-and-Solve and maintains competitive performance with the strong method ARR (Analyzing, Retrieving, and Reasoning).
-Additionally, the effectiveness and generalizability of SWI are solidified on reasoning-intensive question answering (QA) and text summarization benchmarks, where SWI brings consistent improvement to the Baseline generation.
-In text summarization, SWI-generated summaries exhibit greater accuracy, conciseness, and factual correctness, with fewer hallucinations.
-Furthermore, human evaluations verify the coherence, effectiveness, and interpretability of the intent produced by SWI.
-This proof-of-concept study creates a novel avenue for enhancing LLMs’ reasoning abilities with cognitive notions.
-[^1] Appendix A Reproducibility Statement.
-Source code: https://github.com/YuweiYin/SWI
+Extensive experiments on text summarization, multi-task question answering, and mathematical reasoning benchmarks consistently demonstrate the effectiveness and generalizability of Speaking with Intent over direct generation without explicit intent.
+Further analysis corroborates the generalizability of SWI under different experimental settings.
+Moreover, human evaluations verify the coherence, effectiveness, and interpretability of the intent produced by SWI.
+The promising results in enhancing LLMs with explicit intents pave a new avenue for boosting LLMs’ generation and reasoning abilities with cognitive notions.
+[^1] Source code: https://github.com/YuweiYin/SWI
+SWI: Speaking with Intent in Large Language Models
+Yuwei Yin1, EunJeong Hwang1,2, Giuseppe Carenini1 1University of British Columbia; 2Vector Institute for AI {yuweiyin,ejhwang,carenini}@cs.ubc.ca
 1 Introduction
-Large language models (LLMs) (Zhao et al., 2023; Min et al., 2023; Minaee et al., 2024) have revolutionized Natural Language Processing (NLP) with their excellent generative capabilities (Hurst et al., 2024; Anthropic, 2024; Team et al., 2024).
-Many recent benchmarks focus on evaluating complex reasoning in mathematics problems (Cobbe et al., 2021; Hendrycks et al., 2021b; Lightman et al., 2024; Vendrow et al., 2025) and multidisciplinary questions (Hendrycks et al., 2021a; Srivastava et al., 2023; Suzgun et al., 2023; Wang et al., 2024e), presenting significant challenges for LLMs.
-Enhancing LLM reasoning is vital for their ongoing development (Huang & Chang, 2022; Qiao et al., 2023; Patil, 2025).
-Recent advancements have introduced various approaches to improve LLM reasoning (Yu et al., 2024; Cheng et al., 2025; Liu et al., 2025a; Xu et al., 2025).
-Chain-of-Thought (CoT) prompting enhances reasoning (Li et al., 2024b) by providing step-by-step solutions in few-shot exemplars (Wei et al., 2022) or explicitly instructing the model to “think step by step” (Kojima et al., 2022).
-Beyond CoT variants (Yao et al., 2023a; Wang et al., 2023; Yasunaga et al., 2024), Yin & Carenini (2025) propose Analyzing, Retrieving, and Reasoning (ARR), an answer-trigger prompting method that systematically guides LLMs in question answering and outperforms CoT prompting.
-Inspired by ARR (Yin & Carenini, 2025), where intent analysis is identified as its most effective component, we further investigate the powerful potential that intent can unleash in enhancing LLM reasoning.
-Intent, the goal-oriented intention in our mind (Adams, 1986; Mele, 1989; Mele & Moser, 1994), serves as a guiding framework for problem-solving.
-As illustrated in Figure 1(a), human reasoning (System 2 (Kahneman, 2011)) typically follows a structured thinking loop where intent directs problem analysis and logical reasoning.
-Hence, we hypothesize that enabling LLMs to explicitly speak with their own intent—rather than merely analyzing the intent of questions—can replicate this meta-cognitive planning process, thereby improving the reasoning ability and generation quality of LLM.
+Intent, the goal-oriented intention in our mind (Adams, 1986; Mele, 1989; Mele and Moser, 1994), serves as a critical component in communication and a guiding framework for problem-solving.
+As illustrated in Figure 1(a), human thinking (Kahneman, 2011) typically follows a structured loop where intent—a mental state or proactive commitment to perform a specific action or produce a particular outcome—directs problem analysis and logical reasoning, therefore also facilitating communication and interaction.
+Hence, we hypothesize that enabling AI systems to speak with their intent explicitly can replicate this meta-cognitive process, thereby improving their generation quality and reasoning ability.
 Figure 1: SWI Overview.
 (a) The intent, functioning as meta-thought and planning, guides the analysis with reasoning to answer the question.
-(b) A concrete example generated by our method: Speaking with Intent (SWI) in large language models (LLMs).
-(c) The performance improvement on mathematical reasoning and multiple-choice QA tasks brought by SWI.
-This work introduces Speaking with Intent (SWI), requiring LLMs to articulate their own intent as a planning mechanism during generation.
-Due to the autoregressive nature of LLMs (Radford et al., 2019) and the attention mechanism (Vaswani et al., 2017), explicitly stated intent provides high-level guidance for subsequent analysis and reasoning.
-Specifically, we implement SWI using instruction-following LLMs (Ouyang et al., 2022; Rafailov et al., 2023), incorporating system and user prompts for the model (Dubey et al., 2024) to generate intent.
-As shown in Figure 1(b), each analytical step (in blue) in solving a math problem is guided by a preceding intent statement (in orange), which is a freely generated text instead of a predefined class as in traditional intent modeling (Weld et al., 2022).
-To comprehensively evaluate the effectiveness and generalizability of the proposed SWI method, we conduct experiments across three diverse task categories: mathematical reasoning, multiple-choice QA, and text summarization.
-As shown in Figure 1(c), LLM speaking with intent (SWI) consistently outperforms Baseline, which generates responses without explicit intent.
-Furthermore, SWI surpasses answer-trigger prompting methods such as Chain-of-Thought (CoT) (Kojima et al., 2022) and Plan-and-Solve (PS) (Wang et al., 2023).
-Compared with the strong method ARR (Analyzing, Retrieving, and Reasoning) (Yin & Carenini, 2025), SWI maintains competitive performance on mathematical reasoning tasks.
-For text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations (Ji et al., 2023; Li et al., 2024a) in the output.
+(b) The performance improvements brought by SWI on various tasks.
+In recent years, large language models (LLMs) (Zhao et al., 2023; Min et al., 2023; Minaee et al., 2024) have revolutionized Natural Language Processing (NLP) with their excellent generative capabilities (OpenAI, 2024; Anthropic, 2024; Gemini, 2024).
+Enhancing LLMs on various language understanding and logical reasoning tasks (Hendrycks et al., 2021; Suzgun et al., 2023; Lightman et al., 2024; Vendrow et al., 2025) is vital for their ongoing development (Huang and Chang, 2023; Qiao et al., 2023; Patil, 2025).
+This work introduces Speaking with Intent (SWI), requiring LLMs to articulate their own intent as a planning mechanism during generation, which makes LLM intentionality more explicit, in a way reminiscent of the long tradition of intent-driven generation in classic NLG (Grosz and Sidner, 1986; Mann and Thompson, 1988; Moore and Paris, 1993; Reiter and Dale, 2000).
+In essence, we hypothesize that, due to the autoregressive nature of LLMs (Radford et al., 2019) and the attention mechanism (Vaswani et al., 2017), explicitly stated intent provides high-level guidance for subsequent analysis and reasoning.
+For example, when applying SWI to summarization tasks, each analytical step in summarizing an article is effectively guided by a preceding intent statement, which is a piece of free-form text generated by instruction-following LLMs (Ouyang et al., 2022; Rafailov et al., 2023; Llama, 2024) instead of a predefined class as in traditional intent modeling (Weld et al., 2022).
+In this work, we verify the proposed hypothesis by comprehensively evaluating the effectiveness and generalizability of the proposed SWI method.
+Specifically, the experimental results across three diverse task categories (i.e., text summarization, multi-task question answering, and mathematical reasoning) demonstrate that speaking with intent in LLMs consistently outperforms directly generating responses without explicit intent.
+In summarization tasks, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations (Ji et al., 2023; Li et al., 2024a) in the output.
+In math reasoning tasks, SWI surpasses the LLM reasoning method Chain-of-Thought (CoT) (Kojima et al., 2022) and LLM planning method Plan-and-Solve (PS) (Wang et al., 2023), and SWI can work synergistically with these methods to further improve them.
 Additionally, we perform human evaluations to assess the coherence, effectiveness, and interpretability of the intent generated by our SWI method.
 Evaluators largely agree on the quality of the generated intent, particularly for mathematical reasoning tasks.
-The evaluation results confirm that Speaking with Intent in LLMs enhances task performance and output explainability.
-The key contributions of this work are as follows:
-  • 1.
-    This paper introduces Speaking with Intent (SWI) in LLMs, where the generated intent effectively guides problem analysis and logical reasoning, improving performance across various benchmarks.
-The source code will be publicly available.
-  • 2.
-    Extensive experiments across diverse reasoning and generation tasks, including mathematical reasoning, multiple-choice QA, and text summarization, demonstrate that SWI consistently outperforms the Baseline and surpasses established answer-trigger prompting methods such as Chain-of-Thought.
-  • 3.
-    Human evaluations validate the coherence, effectiveness, and interpretability of the intent generated by SWI.
-Rather than intent classification, our evaluation practice provides standards for assessing freely generated intent statements.
-2 Related Work
-LLM Reasoning.
-Recent LLM research increasingly focuses on enhancing reasoning abilities (Sun et al., 2023; Patil, 2025; Cheng et al., 2025; Liu et al., 2025a).
-Chain-of-Thought (CoT) prompting facilitates problem-solving by prompting LLMs to generate intermediate reasoning steps (Kojima et al., 2022; Wei et al., 2022; Li et al., 2024b; Yeo et al., 2025).
-Building on CoT, various reasoning techniques have emerged (Xu et al., 2025).
-Among them, a recent research ARR (Yin & Carenini, 2025) consistently outperforms CoT on multiple QA tasks, where intent analysis of questions is its most effective component.
-Inspired by ARR, we enable LLMs to articulate their intent, using it to guide subsequent analysis and reasoning for improved task performance.
-Results show that Speaking with Intent (SWI) surpasses CoT and ARR on multiple benchmarks.
-As an intent-driven approach, SWI introduces a novel framework for advancing LLM reasoning and cognitive capabilities.
-LLM Planning.
-Planning plays a critical role in advancing artificial intelligence (Wang et al., 2024c), particularly in developing LLM agents in real and complex environments (Qiao et al., 2024; Mialon et al., 2024; Xu et al., 2024; Drouin et al., 2024; Boisvert et al., 2024; Xie et al., 2024a; b).
-Aside from improving the tool-use capabilities (Wang et al., 2024f; Wu et al., 2024; Huang et al., 2024; Wang et al., 2024a; Liu et al., 2025b), recent LLM planning methods enhance reasoning abilities (Huang et al., 2022; Yao et al., 2023b; Hao et al., 2023; Valmeekam et al., 2023; Jiao et al., 2024) by generating problem-solving plans explicitly (Wang et al., 2023; Li & Zhang, 2024; Lyu et al., 2024) or implicitly (Goyal et al., 2024; Wang et al., 2024d; Pfau et al., 2024; Chen et al., 2024).
-In this work, we require LLMs to explicitly speak with their intent in an iterative and dynamic manner, as shown in Figure 1(a), instead of generating all plans first and then solving the task (Wang et al., 2023; Li & Zhang, 2024).
-The existing iterative planning framework (Lyu et al., 2024) focuses on addressing information retrieval issues in knowledge-intensive long-form generation tasks, while our SWI method generates intents that function as strategic planning to guide problem analysis and logical reasoning.
-Intent-related Research.
-Intent Detection (ID) and New Intent Discovery (NID) (Kumar et al., 2022; Liang et al., 2024; Zhang et al., 2024a; Tang et al., 2024; Zhang et al., 2024c; Qian et al., 2024; Yin et al., 2025), which classify utterances into known or novel intent categories, are longstanding challenges in natural language understanding (Larson et al., 2019; Casanueva et al., 2020; Zhang et al., 2021; Weld et al., 2022).
-Typically, these tasks are approached as classification problems (Wang et al., 2024b; Yoon et al., 2024; Zhang et al., 2024b; Sakurai & Miyao, 2024), where models assign sentences to predefined intent classes.
-In contrast, Speaking with Intent (SWI) generates intent as free-form text rather than fixed categories, enhancing flexibility and fluency.
-SWI naturally integrates intent statements as planning into the reasoning process, providing contextual guidance for subsequent analysis.
+The evaluation results confirm that SWI enhances task performance and output explainability.
+The key contributions of this work are as follows: ➊ We introduce Speaking with Intent in LLMs, where the generated intent effectively guides problem analysis, logical reasoning, and text generation, boosting performance across various benchmarks.
+➋ Extensive experiments and analyses across diverse task types and multiple datasets, including text summarization, multi-task QA, and mathematical reasoning, demonstrate the consistent effectiveness and generalizability of SWI.
+➌ Human evaluations validate the coherence, effectiveness, and interpretability of the intent generated by SWI, with our evaluation practice providing standards for assessing freely generated intents.
 2 Speaking with Intent
-This section presents the problem-solving workflow of LLMs and introduces Speaking with Intent (SWI), enabling LLMs to explicitly articulate their intent for planning and reasoning.
+This section presents the problem-solving workflow of LLMs and introduces Speaking with Intent (SWI), enabling LLMs to explicitly articulate their intent during response generation.
 2.1 Problem-solving Workflow using LLMs
 Let $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ be a dataset, where $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n}\}$ is the input information (questions), $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$ is the corresponding references (correct answers), and $n$ is the number of instances in $\mathcal{D}$.
+For text summarization datasets, $X_{i}$ is the source article, and $y_{i}$ is one of the reference summaries.
+For multiple-choice QA datasets, $X_{i}$ contains the question and options, and $y_{i}$ is the answer label such as (A)/(B)/(C) or Yes/No.
 For mathematical reasoning datasets, $X_{i}$ is the math problem, and $y_{i}$ is the correct answer (usually an integer number).
-For multiple-choice QA datasets, $X_{i}$ contains the question and options, and $y_{i}$ is the answer label such as (A)/(B)/(C) or Yes/No.
-For text summarization datasets, $X_{i}$ is the source article, and $y_{i}$ is the reference summary.
 In this work, we employ instruction-following LLMs (aka Chat LLMs) $\mathcal{M}$ for experiments and apply the chat template with the system prompt $P_{s}$ and user prompt $P_{u}$.
@@
 […]
-3.2 Baseline vs. Speaking with Intent
-In this work, we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts $P_{s}$ and restating the SWI requirement in the user prompt $P_{u}$.
-Figure 2 presents the system prompt and user prompt of Baseline generation ($P_{s}^{\text{base}}$ and $P_{u}^{\text{base}}$) and our SWI method ($P_{s}^{\text{swi}}$ and $P_{u}^{\text{swi}}$).
-Figure 2: LLM Prompts. The system and user prompts of Baseline and our SWI method.
+2.2 LLM Speaking with Intent
+SWI is a novel method that brings a significant cognitive concept (i.e., intent) into LLM generation.
+SWI is implemented in a simple and reproducible approach, i.e., we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts $P_{s}$ and restating the SWI requirement in the user prompt $P_{u}$.
+Table 1 presents the SWI instructions for QA and math tasks.
+For summarization tasks, the fourth requirement in the system prompt is At last, clearly and concisely give your final summary starting with "Final Summary:" and the user prompt becomes Speak with intent and summarize the following article.
+{{article}}.
+As showcased in Figure 2 (math) and Figure 3 (summarization), LLMs that speak with intent articulate their intents during the thinking, reasoning and communication process and then provide the final answer based on the analysis.
+System Prompt: You are a helpful assistant who speaks with intent.
+During generation, follow all the requirements below: 1.
+Always explicitly state your own intent before speaking each sentence.
+2.
+Each intent statement should explain the sentence that follows.
+3.
+Your intent must start with the "<INTENT>" tag and end with the "</INTENT>" tag.
+The content within the intent tags must begin with "To" followed by a verb, such as "To accomplish a task." 4.
+At last, clearly and concisely give your final answer starting with "Final Answer:" User Prompt: Speak with intent and answer the following question.
+{{question}}
+Table 1: SWI instructions for QA and math tasks.
+Figure 2: An example of LLMs speaking with intent for mathematical reasoning tasks.
 2.3 Result Evaluation
-Lastly, we extract the final answer (after “Final Answer:” in $\hat{y}_{i}$), denoted as $\tilde{y}_{i}$, and compute the overall performance of $\mathcal{M}$ on the dataset $\mathcal{D}$ by
+We extract the final answer (denoted as $\tilde{y}_{i}$) from the model outputs $\hat{y}_{i}$ and compute the overall performance of $\mathcal{M}$ on the dataset $\mathcal{D}$ by
 $s=\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(y_{i},\tilde{y}_{i}),$ (2)
 where the score function $\mathcal{S}(\cdot,\cdot)$ returns a value in the range of $[0,1]$. Different tasks adopt different score functions to evaluate the model performance.
+To evaluate the quality of summaries, we apply the standard ROUGE (Lin, 2004) as the automatic evaluation metric $\mathcal{S}$ and complement it with a more sophisticated fact-checking analysis as described in Section 4.1.
+For multiple-choice QA tasks, we adopt the Option Selection metric introduced by Yin and Carenini (2025), which evaluates the LLM perplexity of different option concatenations and selects the one with lowest perplexity as the model’s choice.
 For mathematical reasoning tasks, we first extract numbers in $\tilde{y}_{i}$ and apply text normalization to both $\tilde{y}_{i}$ and the reference $y_{i}$, and then conduct exact match to check if the generated answer $\tilde{y}_{i}$ is correct.
-For multiple-choice QA tasks, we adopt the Option Selection metric introduced by Yin & Carenini (2025), which evaluates the LLM perplexity of different option concatenations and selects the one with lowest perplexity as the model’s choice.
-To evaluate the quality of summaries, we apply the standard ROUGE (Lin, 2004) as the automatic evaluation metric $\mathcal{S}$ and complement it with a more sophisticated fact-checking analysis as described in Section 5.2.
-4 Main Experiments on Mathematical Reasoning
-This section presents the main experiments to verify the effectiveness of SWI on mathematical reasoning.
-To ensure reproducibility, we fixed all random seeds, set the generation temperature to zero, and conducted all experiments twice, yielding reproducible results.
-[^2] Please refer to Appendix A for the reproducibility statement and Appendix B for dataset details.
-4.1 Models
-In all experiments, we employ LLaMA3-8B-Instruct (Dubey et al., 2024) as the model $\mathcal{M}$ for generation and evaluation.
+3 Experimental Setup
+This section presents the experimental setup, including Tasks and Datasets (§ 3.1), Model Settings (§ 3.2), and Baseline Settings (§ 3.3).
+3.1 Tasks and Datasets
+The proposed SWI method aims to enhance the generation quality and reasoning ability of LLMs.
+To comprehensively study the effectiveness and generalizability of SWI, we conduct extensive experiments on various challenging benchmarks of three diverse task types: text summarization (Sum), multi-task question answering (QA), and mathematical reasoning (Math).
+The dataset statistics are presented in Table 2.
+Text Summarization (Sum).
+We hypothesize that Speaking with Intent benefits natural language generation tasks like summarization, where the generated intent can guide the model in summarizing the source article point by point in an orderly fashion.
+Hence, we test the effect of SWI on the following text summarization datasets covering different genres: CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017), Extreme summarization (XSum) (Narayan et al., 2018), XL-Sum (Hasan et al., 2021), DialogSum (Chen et al., 2021), and WikiLingua (Ladhak et al., 2020).
+Task Dataset Split Size
+Sum
+CDM Test 11,490
+XSum Test 11,334
+XL-Sum Test 11,535
+DialogSum Test 1,500
+WikiLingua Test 3,000
+QA
+BBH Test 5,511
+MMLU Test 13,842
+MMLU-Pro Test 12,032
+Math
+GSM8K Test 1,319
+GSM8K-P Test 1,209
+MATH500 Test 500
+Table 2: Dataset Statistics.
+Multi-task Question Answering (QA).
+Our SWI method is also evaluated on various multi-task question answering datasets, including BIG-Bench Hard (BBH) (Suzgun et al., 2023), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024b).
+They are all reasoning-intensive benchmarks designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question.
+Here, the hypothesis is that generating its intent explicitly as text improves the system’s question analysis abilities.
+Mathematical Reasoning (Math).
+Beyond multiple-choice QA tasks, we also consider mathematical reasoning benchmarks, where the model is asked to solve the given math problem and present the final answer (numerical values).
+We consider representative and high-quality math benchmarks, including Grade School Math 8K (GSM8K) (Cobbe et al., 2021), GSM8K-Platinum (Vendrow et al., 2025), and MATH500 (Lightman et al., 2024).
+Again, our hypothesis is that speaking with explicit intent improves the model’s reasoning abilities.
+3.2 Model Settings
+Language Models.
+By default, we employ LLaMA3-8B-Instruct (Llama, 2024) as the language model $\mathcal{M}$ for generation and evaluation.
 It is an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters.
-We use the model checkpoint and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020).
-4.2 Datasets
-The proposed SWI method aims to enhance LLM reasoning, so we conduct extensive experiments on various mathematical reasoning benchmarks, ranging from grade-school to competition-level problems.
-In the math reasoning task, the model is asked to solve the given math problem and present the final answer.
-We consider grade-school math problems including GSM8K (Cobbe et al., 2021), GSM8K-Platinum (Vendrow et al., 2025), and MATH500 (Lightman et al., 2024), as well as three competition-level math benchmarks: AMC23, AIME24, and AIME25, which have 40, 30, and 30 math problems, respectively.
-Despite the smaller scale, these math problems are much more challenging than those in GSM8K and MATH500.
-The dataset statistics are presented in Table 1.
-[…]
-Mathematical Reasoning GSM8K (Cobbe et al., 2021) Test 1,319 124
-GSM8K-P (Vendrow et al., 2025) Test 1,209 124
-MATH500 (Lightman et al., 2024) Test 500 134
-[…]
-4.3 Experimental Comparison
-The main comparison is LLM generation with intent (SWI) or without intent (Baseline), and the effectiveness of SWI is verified if the former outperforms the latter.
-For mathematical reasoning and multiple-choice QA tasks, we also compare SWI with answer-trigger prompting methods CoT (Kojima et al., 2022) and ARR (Yin & Carenini, 2025), as well as previous LLM planning method Plan-and-Solve (PS) prompting (Wang et al., 2023).
-CoT aims to elicit LLM reasoning using the answer-trigger prompt $\Phi_{i}^{\text{CoT}}$ as “Let’s think step by step”, while ARR, with a more systematic design, utilizes an enhanced $\Phi_{i}^{\text{ARR}}$ as “Let’s analyze the intent of the question, find relevant information, and answer the question with step-by-step reasoning.” Finally, PS applies the following prompt $\Phi_{i}^{\text{PS}}$ to construct plans before problem-solving: “Let’s first understand the problem and devise a plan to solve the problem.
-Then, let’s carry out the plan and solve the problem step by step.” With answer-trigger prompts $\Phi_{i}$, the generation process (Eq.
+We load the model checkpoint and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020).
+To further assess the generality of SWI, we also evaluate the efficacy of SWI with different LLMs in § 5.3.
+In the fact-checking evaluation for summaries (§ 5.1), we adopt GPT-4o-mini (OpenAI, 2024) to decompose the generated summary and reference summary into two sets of atomic facts.
+Generation Configurations.
+Each experiment session was conducted on a single NVIDIA H100 GPU, and all the models were loaded in a half-precision mode (float16).
+The input sequence is not truncated to avoid losing context information, while we set the maximum number of newly generated tokens to $4096$ during generation.
+Reproducibility Statement.
+To guarantee reproducibility, we fixed the seeds to 42 for all random modules, set the LLM generation temperature to 0 for deterministic generation without sampling, and ran all experiments three times, obtaining reproducible generation outputs and evaluation scores.
+The source code is available on GitHub.
+3.3 Baseline Settings
+The main comparison is LLM generation with intent or without intent, and the effectiveness of SWI is verified if the former outperforms the latter.
+In § 5.2, we also investigate the synergy between SWI and existing LLM reasoning & planning methods, i.e., Chain-of-Thought (CoT) (Kojima et al., 2022) and Plan-and-Solve (PS) (Wang et al., 2023).
+CoT aims to elicit LLM reasoning using the answer-trigger prompt $\Phi_{i}^{\text{CoT}}$ as “Let’s think step by step”, while PS applies the following prompt $\Phi_{i}^{\text{PS}}$ to construct plans before problem-solving: “Let’s first understand the problem and devise a plan to solve the problem.
+Then, let’s carry out the plan and solve the problem step by step.” With such answer-trigger prompts $\Phi_{i}$, the generation process (Eq.
 1) is given by
@@
 where $P_{s}$ is the system prompt, $P_{u}$ is the user prompt, and $X_{i}$ is the task input. $\Theta$ and $\zeta$ are the parameters and hyper-parameters of the LLM $\mathcal{M}$, respectively.
-4.4 Experimental Results
-SWI vs. Baseline.
-As shown in Table 2, SWI improves Baseline (generation without intent) performance by an average of +3.34 points, showing its effectiveness in enhancing LLM’s problem analysis and mathematical reasoning abilities.
-Method GSM8K GSM8K-P MATH500 Competition Avg.
-Baseline 78.47 81.14 36.80 10.56 51.74
-SWI (ours) 81.20 84.20 41.60 13.33 55.08
-Table 2: Mathematical Reasoning Results. The exact matching scores (%) of LLaMA-3-8B-Instruct on multiple mathematical reasoning datasets with or without Speaking with Intent.
-SWI vs. Answer-Trigger Prompting Methods.
-Table 3 presents the results of our SWI method and three answer-trigger prompting methods: CoT, PS, and ARR.
-We observe that SWI significantly outperforms PS, suggesting that the dynamic generation of intents in SWI provides more effective analytical guidance than the static construction of multiple plans at the outset, as in PS.
-Furthermore, compared to the reasoning-eliciting methods CoT and ARR, SWI demonstrates superior performance over CoT and achieves competitive results with ARR.
-Notably, SWI surpasses ARR in solving competition-level mathematics problems, which are more challenging and thus indicate the potential of SWI.
-Method GSM8K GSM8K-P MATH500 Competition Avg.
-CoT (Kojima et al., 2022) 79.30 82.22 42.60 7.78 52.98
-PS (Wang et al., 2023) 73.54 76.43 39.20 8.61 49.45
-ARR (Yin & Carenini, 2025) 81.96 85.36 44.40 10.28 55.50
-SWI (ours) 81.20 84.20 41.60 13.33 55.08
-Table 3: Mathematical Reasoning Results.
-The exact matching scores (%) of LLaMA-3-8B-Instruct on multiple mathematical reasoning datasets using our SWI method or answer-trigger prompting methods.
-The underlined values indicate the second-best scores.
-5 Generalizability of Speaking with Intent
-The main experiments have verified the effectiveness of Speaking with Intent, this section further studies the generalizability of the proposed SWI method on different categories of tasks: multiple-choice QA and text summarization.
-5.1 Multiple-choice Question Answering
-Many challenging benchmarks are designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question.
-5.1.1 QA Datasets
-Our SWI method is evaluated on various QA datasets involving reading comprehension (Liu et al., 2020), commonsense reasoning (Talmor et al., 2019; Sap et al., 2019), world knowledge (Mihaylov et al., 2018), and reasoning-intensive multitask understanding tasks with multiple subtasks covering diverse topics (Clark et al., 2018; Suzgun et al., 2023; Hendrycks et al., 2021a; Wang et al., 2024e).
-The dataset statistics are presented in Table 4.
-[…]
-Reading LogiQA (Liu et al., 2020) Test 651 185
-Commonsense CSQA (Talmor et al., 2019) Valid 1,221 127
-SIQA (Sap et al., 2019) Valid 1,954 117
-World Knowledge OBQA (Mihaylov et al., 2018) Test 500 127
-ARC (Clark et al., 2018) Test 3,548 144
-Multitask Understanding BBH (Suzgun et al., 2023) Test 5,511 203
-MMLU (Hendrycks et al., 2021a) Test 13,842 193
-MMLU-Pro (Wang et al., 2024e) Test 12,032 270
-[…]
-5.1.2 QA Results
-Table 5 shows that our SWI method consistently improves the Baseline (generation without intent) by a large margin, i.e., +10.08 points on average.
-The results demonstrate the efficacy of Speaking with Intent in reasoning-intensive question answering.
-Method Reading Commonsense World Knowledge Multitask Understanding Avg.
-LogiQA CSQA SIQA OBQA ARC BBH MMLU MMLU-Pro
-Baseline 20.74 66.91 29.58 70.60 76.44 56.65 52.40 39.27 51.57
-SWI (ours) 33.03 72.73 46.37 81.60 87.82 64.29 62.19 45.16 61.65
-Table 5: Multiple-choice QA Results. The accuracy scores (%) of LLaMA3-8B-Instruct on various multiple-choice QA datasets with or without Speaking with Intent.
+4 Main Results
+This section presents the experimental results to verify the effectiveness of SWI on diverse generation and reasoning tasks.
+Method Text Summarization (Average ROUGE-1/2/L/Lsum %) Multi-task QA (Accuracy %) Math Reasoning (Accuracy %)
+CDM XSum XL-Sum DialogSum WikiLingua BBH MMLU MMLU-Pro GSM8K GSM8K-P MATH500
+w/o SWI 23.38 11.90 11.29 16.92 15.01 56.65 52.40 39.27 79.08 81.89 38.20
+SWI 23.91 13.90 13.80 19.57 16.53 63.11 59.22 43.72 80.06 82.88 43.00
+Table 3: Main results on text summarization, multi-task QA, and mathematical reasoning tasks.
 4.1 Text Summarization
-The results on math and QA benchmarks have solidified the efficacy of SWI on reasoning-intense tasks.
-Beyond reasoning tasks, we hypothesize that Speaking with Intent benefits natural language generation tasks like summarization by more explicitly analyzing the source document point by point and better planning the generation of the final summary.
-5.2.1 Summarization Datasets
-To verify the hypothesis, we test the effect of SWI on various text summarization datasets.
-CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017), Extreme summarization (XSum) (Narayan et al., 2018), and XL-Sum (Hasan et al., 2021) contain diverse article-summary pairs from news articles.
-SAMSum (Gliwa et al., 2019) and DialogSum (Chen et al., 2021) are dialogue summarization datasets.
-WikiLingua (Ladhak et al., 2020) extracts article and summary pairs from wikiHow.
-The dataset statistics are presented in Table 6.
-[…]
-News Article CDM (Hermann et al., 2015) Test 11,490 920
-XSum (Narayan et al., 2018) Test 11,334 542
-XL-Sum (Hasan et al., 2021) Test 11,535 613
-Dialogue SAMSum (Gliwa et al., 2019) Test 819 209
-DialogSum (Chen et al., 2021) Test 1,500 263
-Wiki Article WikiLingua (Ladhak et al., 2020) Test 3,000 525
-[…]
-5.2.2 Summarization Results
-Automatic Evaluation of Summaries.
+First, we demonstrate that SWI benefits natural language generation tasks like summarization by more explicitly analyzing the source document point by point and better planning the generation of the final summary.
 We evaluate the quality of summaries using the ROUGE score (Lin, 2004), which counts the overlaps of the generated summaries and reference summaries.
 Specifically, we average the ROUGE-1 (unigrams), ROUGE-2 (bigrams), ROUGE-L (longest common subsequences), and ROUGE-LSum (sentence-level ROUGE-L) scores as the final ROUGE score.
-As shown in Table 7, our SWI method consistently surpasses the Baseline on ROUGE scores, confirming its effectiveness in enhancing the quality of text summaries.
-Method News Article Dialogue Wiki Article Avg.
-CDM XSum XL-Sum SAMSum DialogSum WikiLingua
-Baseline 23.38 11.90 11.29 24.14 16.92 15.01 17.11
-SWI (ours) 24.00 13.82 13.50 24.37 20.64 17.86 19.03
-Table 7: Text Summarization Results. The ROUGE scores (%) of the LLaMA3-8B-Instruct model on various text summarization datasets with or without Speaking with Intent (SWI).
-5.2.3 Fact Checking of Summaries
+As shown in Table 3, our SWI method consistently surpasses the direct generation baseline (“w/o SWI”), confirming its effectiveness in enhancing the quality of text summaries.
+4.2 Multi-task Question Answering
+Beyond text understanding and generation, we also consider multi-task QA tasks.
+We test the effect of SWI on three large-scale and challenging benchmarks designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question.
+Table 3 shows that our SWI method consistently improves the direct generation baseline by a large margin.
+The results demonstrate the efficacy of SWI in reasoning-intensive QA tasks.
+4.3 Mathematical Reasoning
+Additionally, we explore the efficacy of SWI on three high-quality math benchmarks.
+Unlike multiple-choice QA in § 4.2, where the model picks an option from the given list, math tasks require LLMs to generate numerical values as the answer.
+As shown in Table 3, SWI consistently improves the model performance over direct generation, showing its effectiveness in enhancing LLM on problem analysis and mathematical reasoning.
+5 Analysis
+Dataset Method Precision Recall F1
+CDM w/o SWI 26.06 76.28 36.37
+SWI 34.22 55.89 37.79
+XSum w/o SWI 11.06 48.38 15.15
+SWI 14.77 37.30 16.29
+XL-Sum w/o SWI 8.96 61.88 13.79
+SWI 12.96 46.72 16.51
+DialogSum w/o SWI 23.99 57.08 29.55
+SWI 34.92 45.19 31.20
+WikiLingua w/o SWI 23.33 65.55 30.63
+SWI 32.40 54.98 35.78
+Table 4: Fact Checking Evaluation of Summaries. We compare the atomic facts drawn from the LLM-generated summaries and the golden references, and compute recall, precision, and F1 scores (%).
+5.1 Fact Checking of Summaries
 LLMs frequently generate hallucinated content (Ji et al., 2023; Li et al., 2024a), which can not be detected by lexical metrics like ROUGE.
-To assess this issue, we adopt a more semantically sophisticated fact-checking metric proposed by Hwang et al.
-(2025), which quantifies factual consistency by calibrating the extent of fabricated statements (low precision) and omitted factual information (low recall).
-Specifically, we use GPT-4o-mini (Hurst et al., 2024) to decompose both generated and reference summaries into atomic fact sets and measure their coverage to quantify factual consistency.
-We evaluate $100$ samples from each summarization dataset using this fact-checking metric, with results presented in Table 8.
-Speaking with Intent (SWI) generates more concise and accurate summaries, improving precision, whereas baseline summaries tend to be more lengthy and verbose, resulting in higher recall scores.
-Overall, SWI consistently outperforms the Baseline generation in terms of F1 score.
-Dataset Method Precision Recall F1 Dataset Method Precision Recall F1
-CDM Baseline 26.06 76.28 36.37 SAMSum Baseline 39.95 75.92 46.90
-SWI 34.22 55.89 37.79 SWI 47.82 69.88 50.88
-XSum Baseline 11.06 48.38 15.15 DialogSum Baseline 23.99 57.08 29.55
-SWI 14.77 37.30 16.29 SWI 34.92 45.19 31.20
-XL-Sum Baseline 8.96 61.88 13.79 WikiLingua Baseline 23.33 65.55 30.63
-SWI 12.96 46.72 16.51 SWI 32.40 54.98 35.78
-Table 8: Summarization evaluation results. We compare the atomic facts drawn from the generated and reference summaries, and compute recall, precision, and F1 scores (%).
-5.2.4 Case Study
-To provide more insights into the benefits of Speaking with Intent, we conduct a case study on the CDM summarization dataset (Hermann et al., 2015; See et al., 2017).
+To assess this issue, we adopt a more semantically sophisticated fact-checking metric (Hwang et al., 2025), which quantifies factual consistency by calibrating the extent of fabricated statements (low precision) and omitted factual information (low recall).
+Specifically, we use GPT-4o-mini (OpenAI, 2024) to decompose both generated and reference summaries into atomic fact sets and measure their coverage to quantify factual consistency.
+We evaluate $100$ random samples from each summarization dataset using this fact-checking metric, with results presented in Table 4.
+Directly generated summaries (“w/o SWI”) tend to be more lengthy and verbose, resulting in higher recall scores.
+In contrast, SWI-generated summaries exhibit greater accuracy, conciseness, and factual correctness, with fewer hallucinations.
+Overall, SWI consistently outperforms the direct generation baseline in terms of F1 score.
+Method GSM8K GSM8K-P MATH500 Avg.
+CoT w/o SWI 77.86 80.07 42.00 66.64
+SWI 80.21 82.55 42.80 68.52
+PS w/o SWI 72.56 75.35 40.00 62.64
+SWI 79.45 82.54 41.40 67.80
+Table 5: LLM Reasoning & Planning with SWI. When additional LLM reasoning (CoT) and planning (PS) methods are adopted, the exact matching scores (%) on multiple math datasets with or without SWI.
+Method News Article Dialogue Wiki Article
+CDM XSum XL-Sum DialogSum WikiLingua
+CoT w/o SWI 23.17 11.54 11.11 15.77 14.44
+SWI 24.25 13.86 13.73 19.49 16.88
+PS w/o SWI 24.12 12.21 11.91 17.92 15.86
+SWI 24.43 12.46 12.28 18.95 16.76
+Table 6: LLM Reasoning & Planning with SWI. When additional LLM reasoning (CoT) and planning (PS) methods are adopted, the ROUGE scores (%) on multiple summarization datasets with or without SWI.
+5.2 Synergy with Other Methods
+In recent years, various methods have been proposed to boost the reasoning and planning abilities of LLMs.
+Since our SWI method is orthogonal to previous work, it is necessary to compare the performance and study the synergy between SWI and them.
+As mentioned in § 3.3, we adopt representative LLM reasoning method Chain-of-Thought (CoT) (Kojima et al., 2022) and LLM planning method Plan-and-Solve (PS) (Wang et al., 2023).
+When additional reasoning (CoT) and planning (PS) methods are adopted, Table 5 and Table 6 present the LLM performance on multiple math and summarization tasks, respectively.
+Comparing the results in Table 3 and Table 5, our SWI method (➁ in Table 3) beats CoT (➀ in Table 5) and PS (➂ in Table 5).
+Moreover, the combination of CoT+SWI (➁ in Tables 5 and 6) boosts the CoT method (i.e., ➁>>➀), and the synergy of PS+SWI (➃ in Tables 5 and 6) also improves the PS-alone performance (i.e., ➃>>➂).
+These results verify that SWI works synergistically with existing LLM reasoning & planning methods.
+5.3 Generalizability to Different LLMs
+To further validate the generalizability of our SWI method, we evaluate its effect on different sizes and types of LLMs.
+Aside from the results of LLaMA3-8B (Table 3), we present the results on multiple mathematical reasoning tasks using LLaMA3-3B (with 3B parameters) and LLaMA3-8B-R1, which is fine-tuned using reasoning data distilled from DeepSeek R1 (Guo et al., 2025).
+As observed in Table 7, our SWI method brings consistent improvements over the direct generation baseline, verifying the effectiveness of SWI when applied to models of different model sizes and LLM types (i.e., chat and reasoning models).
+Model Method GSM8K GSM8K-P MATH500 Avg.
+LLaMA3-3B w/o SWI 45.64 46.82 27.20 39.89
+SWI 65.05 67.58 32.80 55.14
+LLaMA3-8B-R1 w/o SWI 68.08 70.72 56.40 65.07
+SWI 75.44 79.24 57.00 70.56
+Table 7: Generalizability of SWI to different LLMs. When different sizes and types of LLMs are adopted, the exact matching scores (%) on multiple mathematical reasoning datasets with or without SWI.
+Method News Article Dialogue Wiki Article
+CDM XSum XL-Sum DialogSum WikiLingua
+w/o SWI 23.38 11.90 11.29 16.92 15.01
+SWI (V0) 23.91 13.90 13.80 19.57 16.53
+SWI (V1) 24.27 14.12 14.10 19.43 17.34
+SWI (V2) 24.04 14.69 14.66 19.24 17.06
+SWI (V3) 24.17 13.83 14.02 18.82 16.10
+Table 8: Results of Different SWI Prompt Variants.
+When different paraphrases of SWI prompts are adopted, the ROUGE scores (%) on various text summarization datasets with or without Speaking with Intent (SWI).
+5.4 SWI Prompt Variants
+In this work, we implement SWI in a straightforward prompting way for simplicity and reproducibility.
+To demonstrate that SWI works effectively irrespective of specific prompt design, we conduct experiments on different SWI prompt variants.
+Specifically, the original SWI prompt (“V0”) is paraphrased into three different versions (details in the appendix) by GPT-4o (OpenAI, 2024).
+Table 8 shows the performance of each SWI variants over multiple summarization tasks.
+As illustrated, the proposed SWI method, regardless of its prompt implementation, maintains a consistent advantage over the direct generation baseline, substantiating that SWI is an effective and general framework that brings intent—a key cognitive concept for reasoning and communication—into LLM generation.
+Task Dataset # Input Tokens # Output Tokens
+w/o SWI SWI $\Delta$ w/o SWI SWI $\Delta$
+Sum CDM 920 1028 +108 166 434 +161%
+XSum 542 650 +108 135 374 +177%
+XL-Sum 613 721 +108 150 405 +170%
+DialogSum 263 371 +108 77 238 +209%
+WikiLingua 525 633 +108 145 386 +166%
+QA BBH 203 313 +110 96 244 +154%
+MMLU 193 303 +110 63 229 +263%
+MMLU-Pro 270 380 +110 493 703 +43%
+Math GSM8K 123 233 +110 211 280 +33%
+GSM8K-P 123 233 +110 213 273 +28%
+MATH500 133 243 +110 810 922 +14%
+Table 9: Efficiency Study. The number of input and output tokens with or without SWI.
+5.5 Efficiency Study
+To study the efficiency cost of requiring LLMs to speak with intent, we present the number of input and output tokens with or without SWI in Table 9.
+SWI adds about $110$ extra input tokens in the system prompt (as in Table 1) for additional instructions, and the extra output tokens are not more than 300 tokens (percentages in Table 9) on average compared to direct generation.
+Hence, this cost should be traded off against the substantially improved performance, transparency, and interpretability of SWI.
+Figure 3: Case Study. The source article, reference summary, and SWI output of a text summarization example.
+5.6 Case Study
+In addition to the math example in Figure 2, we conduct a case study on the summarization task to provide more insights into the benefits of SWI.
 Figure 3 showcases the SWI outputs, where the generated intent is well formulated and articulated, guiding a point-by-point summarization process that leads to a final summary that is accurate, concise, and abstractive, while effectively capturing the key information of the source article.
-This probably leads to the high Precision and F1 scores of SWI in Table 8, suggesting the validity of Speaking with Intent in text generation tasks.
-Figure 3: Case Study. Task input and SWI output on text summarization task.
-6 Intent Quality Evaluation
-Although we have shown that Speaking with Intent boosts performance across a broad range of tasks, verifying the quality of generated intents is also significant.
+This probably leads to the high Precision and F1 scores of SWI in Table 4, suggesting the validity of SWI in text generation tasks.
+Similarly, SWI enables LLM to have progressive planning when solving math problems.
+This is a critical ability when dealing with complex problems that require the divide-and-conquer strategy.
+Task Dataset Total Unique Per Ins.
+Sum CDM 53,086 86 4.6
+XSum 50,697 92 4.5
+XL-Sum 56,100 84 4.9
+DialogSum 4,739 96 3.2
+WikiLingua 12,463 91 4.2
+QA BBH 7,425 49 1.3
+MMLU 19,380 106 1.4
+MMLU-Pro 21,916 122 1.8
+Math GSM8K 5,237 45 4.0
+GSM8K-P 4,719 43 3.9
+MATH500 3,619 84 7.2
+Table 10: Intent Statistics. The number of total and unique intents (verbs) of each dataset, as well as the average number of intents per instance (“Per Ins.”).
+5.7 Intent Statistics
+To further analyze the pattern and variability of the intents generated by our SWI method, we present the intent statistics across different tasks.
+Specifically, we extract and count the verbs in the generated intent statements, which are required to be in the format “To do something.”.
+Table 10 shows the number of total verbs (“Total”) and unique verbs (“Unique”) of each dataset, as well as the average number of intents per instance (“Per Ins.”).
+We observe that the number of unique intents in the summarization task is larger, indicating that summarizing documents demands a higher intent variability.
+In addition, summarization and math tasks generally have more intents per instance than QA, likely due to their longer outputs, as also observed in the efficiency study (Table 9).
+Among the three math datasets, MATH500 is relatively harder, as the model performance is lower in Table 3.
+Thus, solving the problems in MATH500 requires more thinking steps and longer reasoning chains, which is consistent with the observation that its number of unique intent verbs and the average number of intents per instance are larger than GSM8K.
+Figure 4: The semantic distribution of intents across different task types and among all English verbs.
+Furthermore, we investigate the intents across different task types and the distribution of intent verbs among all English verbs.
+First, we feed the same model that generates the outputs (i.e., LLaMA3-8B) with each verb and extract the last-layer hidden states, which indicate how the generator perceives and utilizes the intents.
+Then, t-SNE (van der Maaten and Hinton, 2008) is applied to visualize the semantic representations of each unique intent verb and all 11,531 English verbs drawn from WordNet (Miller, 1992).
+We observe from Figure 4(a) that all three tasks involve diverse intents, indicating the need for versatile intent skill sets when performing different tasks.
+Figure 4(b) illustrates that intent verbs mainly lie in a certain cluster, showing their specialty from common English verbs.
+In addition, Figure 5 presents the top 10 common intent verbs in summarization, QA, and math tasks, demonstrating the nuance of intents used in different tasks.
+Figure 5: Top 10 common intent verbs of (a) summarization, (b) question answering, and (c) math tasks.
+5.8 Intent Quality Evaluation
+Although we have shown that SWI boosts performance across a broad range of tasks, verifying the quality of generated intents is also significant.
 Thus, we hire human evaluators to assess the quality of generated intent across three criteria: coherence, effectiveness, and interpretability.
 Coherence measures how well the intent guides analysis and reasoning, effectiveness evaluates its contribution to problem-solving, and interpretability assesses its role in enhancing user understanding of the generated content.
-[^3] Please refer to Appendix C for more details on human evaluation.
 For each instance, human evaluators are provided with evaluation instructions, task input (e.g., the math problem, question with options, or source article), SWI-generated output, and assessment check boxes.
@@
   • Coherence: In general, does the analysis align coherently with the intent statements?
-  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for solving the problem?
+  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for performing the task?
   • Interpretability: Do you think providing the intent can help you better understand the reasoning process than not providing it?
+Task Dataset Coherence Effectiveness Interpretability
+Score Agree Score Agree Score Agree
+Summarization CDM 2.83 80% 2.77 70% 2.83 75%
+XSum 2.70 70% 2.60 65% 2.57 65%
+Math Reasoning GSM8K 2.90 85% 2.97 95% 2.97 95%
+MATH500 2.87 80% 2.87 80% 2.83 80%
+Multi-task QA BBH 2.37 55% 2.37 50% 2.33 45%
+MMLU 2.67 75% 2.53 55% 2.37 45%
+Table 11: Intent Quality Evaluation by Humans. The score ranges from 1 (Bad) to 3 (Good).
 Evaluation scores range from 1 (Bad), 2 (Fair), to 3 (Good).
 Agreement ratios are calculated as follows: 1 if all three evaluators agree, 0.5 if two agree, and 0 if all scores differ.
-As shown in Table 9, human evaluation scores for all aspects across datasets exceed 2, indicating that the generated intent is generally well-regarded.
-Notably, math reasoning tasks achieve scores near 3 with agreement ratios approaching 100%, demonstrating that SWI-generated intent is particularly coherent, effective, and interpretable in mathematical problem-solving.
-The relatively low scores observed in QA tasks may be attributed to the lack of multi-step guidance: the number of generated intents in QA tasks is often 1 or 2, while that in math and summarization is usually at least three.
-This finding indicates the advantages of multi-round iterative intents, although SWI can boost task performance even with few intents generated.
-Task Dataset Coherence Effectiveness Interpretability
-Score Agree Score Agree Score Agree
-Math Reasoning GSM8K 2.90 85% 2.97 95% 2.97 95%
-MATH500 2.87 80% 2.87 80% 2.83 80%
-Multiple-choice QA BBH 2.37 55% 2.37 50% 2.33 45%
-MMLU 2.67 75% 2.53 55% 2.37 45%
-Summarization CDM 2.83 80% 2.77 70% 2.83 75%
-XSum 2.70 70% 2.60 65% 2.57 65%
-Table 9: Intent Quality Evaluation by Humans. The score ranges from 1 (Bad) to 3 (Good).
+As shown in Table 11, human evaluation scores for all aspects across datasets exceed 2.3, indicating that the generated intent is generally well-regarded.
+Notably, we observe very strong assessment scores (near 3) with substantial agreement (approaching 100%) for both summarization and math tasks, demonstrating that SWI-generated intent is particularly coherent, effective, and interpretable.
+The relatively low (but still fairly good) scores observed in QA tasks may be attributed to the lack of multi-step guidance: as presented in Table 10, the average number of intents per instance in QA tasks is often 1 or 2, which is much lower than that in summarization and math tasks.
+This finding indicates the advantages of multi-round iterative intents, with SWI being able to boost task performance even with a few intents generated.
+6 Related Work
+Intent in NLG and LLMs.
+Since the seminal work (Grosz and Sidner, 1986), intent has played a critical role in NLG.
+In the classical NLG pipeline (Reiter and Dale, 1997, 2000), content determination and document planning are modeled as a process of communicative goals decomposition and ordering, where the resulting planned communicative acts the NLG system wants to achieve are its intentions.
+Most approaches followed this framework by implementing a computational model of the Rhetorical Structure Theory (RST) (Mann and Thompson, 1988), a discourse theory that explains how parts of a multi-sentential text relate to each other functionally, i.e., how each piece serves a communicative purpose relative to the whole.
+For instance, RST has been applied in NLG to many genres, ranging from handling explanation dialogues (Moore and Paris, 1993) to generating persuasive evaluative arguments (Carenini and Moore, 2006).
+In all these applications, the text was planned before being generated and intentions were explicit.
+In contrast, in modern LLM-driven generation, intent is typically implicit.
+In other words, what we see is only the generated text, with no access to the underlying communicative goals and corresponding intentions.
+In this respect, SWI can be seen as the first attempt to make LLM intentionality more explicit, bridging the gap between classic NLG and LLMs, arguably boosting LLMs’ controllability and interpretability.
+LLM Reasoning.
+Early LLMs were rather poor at reasoning (Radford et al., 2018, 2019), and scaling pre-training was shown not to be a feasible solution for improving reasoning (Chu et al., 2025).
+Instead, Chain-of-Thought (CoT) prompting (Kojima et al., 2022; Wei et al., 2022) demonstrated that by modifying the prompt, LLMs can elicit a beneficial step-by-step reasoning process at test time without additional training (Li et al., 2024b; Yeo et al., 2025; Zhang et al., 2025).
+Building on CoT, various reasoning techniques have emerged (Xu et al., 2025).
+Among them, a recent research ARR (Yin and Carenini, 2025) consistently outperforms CoT on multiple QA tasks, where analyzing the intent of questions is its most effective component.
+Different from ARR, we enable LLMs to articulate their intent, using it to guide subsequent analysis and reasoning for improved task performance.
+Inspired by the success of CoT and similar prompting techniques, very recent research is increasingly focusing on enhancing LLMs reasoning abilities by explicitly training them for reasoning using reinforcement learning (RL) (Sutton and Barto, 2018; Shao et al., 2024).
+Intriguingly, the success of SWI demonstrated in this paper could spur a similar explosion of research on training LLMs with RL (Wang et al., 2025; Setlur et al., 2025) to better analyze and formulate intentions.
+Intent-related Research.
+Intent Detection (ID) and New Intent Discovery (NID) (Kumar et al., 2022; Liang et al., 2024; Zhang et al., 2024a; Tang et al., 2024; Zhang et al., 2024c; Qian et al., 2024; Yin et al., 2025), which classify utterances into known or novel intent categories, are longstanding challenges in natural language understanding (Larson et al., 2019; Casanueva et al., 2020; Zhang et al., 2021; Weld et al., 2022).
+Typically, these tasks are approached as classification problems (Wang et al., 2024a; Yoon et al., 2024; Zhang et al., 2024b; Sakurai and Miyao, 2024), where models assign sentences to predefined intent classes.
+On the contrary, our SWI method generates intent as free-form text rather than fixed categories, enhancing flexibility and fluency.
+SWI naturally integrates intent statements as planning into the reasoning process, providing contextual guidance for subsequent analysis.
 7 Conclusion
-In this work, we introduce Speaking with Intent (SWI) in LLMs, where the generated intent (as high-level planning) guides subsequent analysis, improving the reasoning and generation abilities.
-Extensive experiments across mathematical reasoning, multiple-choice QA, and text summarization benchmarks consistently show the benefits of SWI over the Baseline method that generates without explicit intent.
-In addition, SWI outperforms answer-trigger prompting methods like Chain-of-Thought on mathematical reasoning benchmarks.
+In this work, we introduce Speaking with Intent (SWI) in LLMs, where the generated intent (as high-level planning) guides subsequent analysis, improving the generation and reasoning abilities.
+Extensive experiments across text summarization, multi-task QA, and mathematical reasoning benchmarks consistently show the benefits of speaking with explicit intent over the direct generation baseline.
 In text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations.
-Furthermore, human evaluations solidify the coherence, effectiveness, and interpretability of LLM-generated intent.
-This study opens a new avenue for enhancing LLM reasoning abilities and beyond.
-Impact Statement.
+In addition, SWI outperforms existing LLM reasoning and planing methods and works synergistically with them.
+Further analysis substantiates the generalizability of SWI when applied to different settings.
+Moreover, human evaluations solidify the coherence, effectiveness, and interpretability of LLM-generated intent.
+Overall, this study opens a new avenue for enhancing LLM generation and reasoning abilities.
+Ethics & Impact Statement
+This work does not raise ethical issues, and we would like to mention the impact of SWI.
 As intent is a fundamental aspect of natural language processing, empowering, eliciting, and enhancing the intent understanding and generation abilities can potentially drive AI systems (including multimodal models) to the next level.
@@
 Acknowledgments
+Nous remercions le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) de son soutien.
 We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), Vector Institute for AI, and Accelerate Foundation Models Research Program Award from Microsoft.
@@
 We would also like to thank Vered Shwartz for constructive feedback on conducting evaluations.
-Appendix A Reproducibility Statement
-A.1 LLM Generation
-For all experiments in § 4 and § 5, we employ LLaMA3-8B-Instruct (Dubey et al., 2024), an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters, and use the model checkpoint and tokenizer [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct provided by Hugging Face Transformers (Wolf et al., 2020).
-Only in the fact-checking evaluation for summaries (§ 5.2.3), we adopt the proprietary model GPT-4o-mini (Hurst et al., 2024) to decompose the generated summary and reference summary into two sets of atomic facts.
-For each running session, the experiments are conducted on a single NVIDIA V100 GPU with 32GB memory.
-To avoid out-of-memory issues, all the models are loaded in a half-precision (float16) mode, and the generation batch size is $1$.
-The input sequence is not truncated since we do not want to lose the context information, but we set the maximum number of newly generated tokens as $4096$ during generation.
-To guarantee reproducibility, we fixed the seeds as 42 for all random modules, set the LLM generation temperature as 0 for deterministic generation without sampling, and ran all experiments twice, obtaining reproducible generation results and evaluation scores.
-The source code is available on GitHub: https://github.com/YuweiYin/SWI
-A.2 Experimental Cost
-In the generation stage, the total computational cost is approximately 1,500 GPU hours on NVIDIA V100 clusters (about 58 days).
-In the evaluation stage, only the option selection process for multiple-choice QA tasks requires GPU usage (Yin & Carenini, 2025), and the overall running time is roughly 100 hours on V100 clusters.
-In the fact-checking experiments (§ 5.2.3), the expense for GPT-4o-mini API calls is below US\$10.
-Appendix B Tasks and Datasets
-To comprehensively study the effectiveness and generalizability of the proposed SWI method, we consider three different categories of tasks in our experiments: mathematical reasoning, multiple-choice QA, and text summarization.
-Table 10 presents the dataset sources.
-Please note that the URLs may be subject to change by the dataset providers.
-Task Dataset URL
-Mathematical Reasoning GSM8K (Cobbe et al., 2021) Link
+Appendix A Experiment Details
+A.1 Dataset Details
+All datasets used in this work are loaded from Hugging Face datasets. Table 12 lists the URL link of each dataset. Please note that the URLs may be subject to change by the dataset providers.
+Task Dataset Source
+Sum
+CDM (Hermann et al., 2015) Link
+XSum (Narayan et al., 2018) Link
+XL-Sum (Hasan et al., 2021) Link
+DialogSum (Chen et al., 2021) Link
+WikiLingua (Ladhak et al., 2020) Link
+QA
+BBH (Suzgun et al., 2023) Link
+MMLU (Hendrycks et al., 2021) Link
+MMLU-Pro (Wang et al., 2024b) Link
+Math
+GSM8K (Cobbe et al., 2021) Link
 GSM8K-P (Vendrow et al., 2025) Link
 MATH500 (Lightman et al., 2024) Link
-AMC23 (American Mathematics Competitions) Link
-AIME24 (American Invitational Math Examination) Link
-AIME25 (American Invitational Math Examination) Link
-Multiple-choice QA LogiQA (Liu et al., 2020) Link
-CSQA (Talmor et al., 2019) Link
-SIQA (Sap et al., 2019) Link
-OBQA (Mihaylov et al., 2018) Link
-ARC (Clark et al., 2018) Link
-BBH (Suzgun et al., 2023) Link
-MMLU (Hendrycks et al., 2021a) Link
-MMLU-Pro (Wang et al., 2024e) Link
-Text Summarization CDM (Hermann et al., 2015) Link
-XSum (Narayan et al., 2018) Link
-XL-Sum (Hasan et al., 2021) Link
-SAMSum (Gliwa et al., 2019) Link
-DialogSum (Chen et al., 2021) Link
-WikiLingua (Ladhak et al., 2020) Link
 Table 12: Dataset Sources.
-B.1 Mathematical Reasoning
-In the mathematical reasoning task, the model is asked to solve the given math problem and present the final answer. We consider the following math datasets.
-GSM8K.
-Grade School Math 8K (GSM8K) (Cobbe et al., 2021) is a dataset of 8.5K high-quality linguistically diverse grade school math word problems.
-The dataset was created to support the task of question answering on basic mathematical problems that require multi-step reasoning.
-GSM8K-P.
-GSM8K-Platinum (Vendrow et al., 2025) is a revised version of the full test set of GSM8K (Cobbe et al., 2021).
-It revises the labels of mislabeled examples and removes any question that is determined to be poorly written (mostly due to ambiguity in the problem statement), providing a more accurate assessment of mathematical reasoning capabilities.
-MATH500.
-MATH500 (Lightman et al., 2024) consists of 500 questions from the MATH test set (Hendrycks et al., 2021b).
-The distribution of difficulty levels and subjects in MATH500 is representative of the MATH test set as a whole.
-Competition.
-We consider three competition-level math benchmarks: AMC23, AIME24, and AIME25, drawn from American Mathematics Competitions and American Invitational Mathematics Examination.
-AMC23, AIME24, and AIME25 have 40, 30, and 30 math problems, respectively.
-Despite the smaller scale, these math problems are much more challenging than those in GSM8K and MATH500.
-B.2 Multiple-choice QA
-Since many challenging benchmarks are designed as multiple-choice question-answering tasks, our SWI method is also evaluated on various reasoning-intense QA datasets, where the model is asked to select the most appropriate one from the given options to answer the question.
-LogiQA.
-LogiQA Liu et al. (2020) is a reading comprehension dataset that requires the model to have logical reasoning for question-answering.
-CSQA.
-CommonsenseQA (CSQA) (Talmor et al., 2019) examines the model on commonsense question-answering problems constructed using information from ConceptNet Speer et al. (2017).
-SIQA.
-SocialIQA (SIQA) (Sap et al., 2019) is a large-scale QA benchmark for commonsense reasoning about social situations, which probes emotional and social intelligence in everyday situations.
-OBQA.
-OpenBookQA (OBQA) (Mihaylov et al., 2018) asks the model to answer the question based on the given elementary-level science facts and broad commonsense knowledge.
-ARC.
-AI2 Reasoning Challenge (ARC) (Clark et al., 2018) contains grade-school science questions.
-It is divided into a Challenge and an Easy set, where the former contains only questions answered incorrectly by both a retrieval-based algorithm and a word co-occurrence algorithm.
-BBH.
-BIG-Bench Hard (BBH) (Suzgun et al., 2023) is a suite of challenging tasks filtered from BIG-Bench (Srivastava et al., 2023).
-Solving these problems often requires multi-step reasoning.
-In this work, $4$ (out of $27$) subtasks in BBH (word_sorting, object_counting, dyck_languages, and multistep_arithmetic_two) are discarded as they are not multiple-choice QA tasks.
-MMLU.
-MMLU (Hendrycks et al., 2021a) comprehensively measures the multitask language understanding abilities on $57$ subtasks including elementary mathematics, history, computer science, and more.
-MMLU-Pro.
-MMLU-Pro (Wang et al., 2024e) extends the mostly knowledge-driven MMLU benchmark by integrating more challenging, reasoning-focused questions and expanding the choice set from four to ten options.
-B.3 Text Summarization
-Beyond math reasoning and multiple-choice QA, we hypothesize that Speaking with Intent benefits natural language generation tasks like summarization, where the generated intent can guide the model in summarizing the source article point by point.
-Hence, we test the effect of SWI on the following text summarization datasets (English only).
-CNN/DailyMail.
-CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017) contains over 300K unique news articles written by journalists at CNN and the Daily Mail.
-XSum.
-Extreme summarization (XSum) (Narayan et al., 2018) is a real-world, large-scale dataset consisting of online articles from the British Broadcasting Corporation (BBC).
-XL-Sum.
-XL-Sum (Hasan et al., 2021) is a comprehensive and diverse dataset comprising 1.35 million professionally annotated article-summary pairs from BBC News.
-The dataset covers 45 languages and we only use the English subset.
-SAMSum.
-SAMSum (Gliwa et al., 2019) contains about 16K messenger-like conversations with summaries.
-Linguists were asked to create conversations similar to those they write on a daily basis, reflecting the proportion of topics of their real-life messenger conversations.
-DialogSum.
-DialogSum (Chen et al., 2021) is a large-scale dialogue summarization dataset, consisting of 13,460 dialogues with corresponding manually labeled summaries and topics.
-WikiLingua.
-WikiLingua (Ladhak et al., 2020) is a large-scale, multilingual dataset for evaluating cross-lingual abstractive summarization systems.
-It extracts article and summary pairs in 18 languages from wikiHow, and we only use the English subset.
-Appendix C Human Evaluation Details
+A.2 Model Details
+As mentioned in § 3.2, we mainly employ LLaMA3-8B-Instruct (Llama, 2024), an instruction-following LLM with 8 billion model parameters, for most experiments.
+In generalizability experiments (§ 5.3), we also explore LLMs of different sizes and types.
+Table 13 presents the URL link of each model and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020).
+Model Source
+LLaMA3-3B (Llama, 2024) Link
+LLaMA3-8B (Llama, 2024) Link
+LLaMA3-8B-R1 (Guo et al., 2025) Link
+Table 13: Model Sources.
+A.3 SWI Prompt Variants
+As mentioned in § 2.2, we implement SWI in a straightforward prompting way for simplicity and reproducibility, i.e., we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts and restating the SWI requirement in the user prompt.
+In addition, we paraphrase the SWI prompt into three different versions (§ 5.4) to demonstrate that our SWI method maintains effectiveness irrespective of the specific prompt formulation.
+Here, we present the prompt variants in Table 14.
+Type & Version Prompt Text
+System Prompt
+V0 (default)
+You are a helpful assistant who speaks with intent. You are good at summarizing
+documents and the summary must start with "Final Summary:"
+During generation, follow all the requirements below:
+1. Always explicitly state your own intent before speaking each sentence.
+2. Each intent statement should explain the sentence that follows.
+3. Your intent must start with the "<INTENT>" tag and end with the "</INTENT>" tag.
+The content within the intent tags must begin with
+"To" followed by a verb, such as "To accomplish a task."
+4. At last, clearly and concisely give your final summary starting with "Final Summary:"
+System Prompt
+V1
+You are a purposeful assistant skilled in document summarization who speaks with intent.
+Your final response must begin with "Final Summary:"
+While generating responses, adhere strictly to these instructions:
+1. Before every sentence, clearly state your intent using an explanation.
+2. Each intention should directly clarify the sentence that follows.
+3. Use the tags "<INTENT>" and "</INTENT>" to wrap each intent statement.
+Each statement inside the intent tags must begin with "To" and a verb,
+for example, "To describe the process."
+4. Conclude with a clear and concise final summary that begins with "Final Summary:"
+System Prompt
+V2
+You are a helpful assistant who is skilled in text summarization and always communicates
+with deliberate intent. Ensure your final output starts with "Final Summary:"
+Comply with the following instructions during your response:
+1. Begin each sentence with a description of your intent.
+2. The intent must directly relate to and explain the sentence that comes after it.
+3. Surround each intent with the tags "<INTENT>" and "</INTENT>". Each intent
+statement enclosed by the tags should start with the word "To" and an action verb,
+like "To explain the reasoning."
+4. Finish with a succinct summary, introduced by "Final Summary:"
+System Prompt
+V3
+You are a precise and helpful assistant proficient in text summarization, who always
+speaks with deliberate intent. Your final response must begin with "Final Summary:"
+While producing your response, follow these guidelines:
+1. Before each sentence, declare your intent explicitly.
+2. Ensure each intent explains the sentence that immediately follows.
+3. Wrap every intent declaration with "<INTENT>" and "</INTENT>" tags.
+Make sure that every intent statement within the tags begins with "To" and an action verb,
+for example, "To justify the choice."
+4. Conclude your response with a clearly stated final summary prefaced by "Final Summary:"
+User Prompt
+All Versions
+Speak with intent and summarize the following document.
+{{article}}
+Table 14: SWI Prompt Variants.
+Appendix B Human Evaluation Details
 Participant Requirements.
-We hire human evaluators from the cloud-sourcing platform CloudResearch [^5] https://www.cloudresearch.com/ to conduct human evaluation on the quality of the generated intent: coherence, effectiveness, and interpretability.
+We hire human evaluators from the cloud-sourcing platform CloudResearch to conduct human evaluation on the quality of the generated intent: coherence, effectiveness, and interpretability.
 To ensure the annotation quality, we apply several requirements to select qualified human evaluators, as shown in Table 15.
@@
 Evaluation Tasks.
-For each task category, we select two datasets: GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2024) for mathematical reasoning, BBH (Suzgun et al., 2023) and MMLU (Hendrycks et al., 2021a) for multiple-choice QA, and CDM (Hermann et al., 2015; See et al., 2017) and XSum (Narayan et al., 2018) for text summarization.
+For each task category, we select two datasets: CDM (Hermann et al., 2015; See et al., 2017) and XSum (Narayan et al., 2018) for text summarization, BBH (Suzgun et al., 2023) and MMLU (Hendrycks et al., 2021) for multi-task multiple-choice QA, and GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2024) for mathematical reasoning.
 We randomly sample 12 instances per dataset and divide them into two batches of six.
@@
   • Coherence: In general, does the analysis align coherently with the intent statements?
-  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for solving the problem?
+  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for performing the task?
   • Interpretability: Do you think providing the intent can help you better understand the reasoning process than not providing it?